### v2

# Hilbert-Kunz multiplicity of fiber products and idealization rings

Abstract.
The main purpose of this paper is to provide formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings. We calculate the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$, with $R$, $S$ and $T$ being Noetherian local rings with common characteristic and residue field. Inspired by a work of Strazzanti and Armengou we calculate the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module. We explore certain examples and structural consequences and obtain some bounds for the Hilbert-Kunz multiplicity.
Key words and phrases:
Hilbert-Kunz, multiplicity, fiber product, idealization
2020 Mathematics Subject Classification:
13D40, 13H15, 13H05
Corresponding author: P. D. C. Martins
The first author was supported by grant 2019/21181-0, São Paulo Research Foundation (FAPESP). The second author was supported by grant 2022/12114-0, São Paulo Research Foundation (FAPESP)
1. Introduction
Throughout this paper, $(R,\mathfrak{m},k)$ will denote a Noetherian local ring of prime characteristic $p>0$ and dimension $d:=\dim R>0$. It’s worth noting that in this context, the residue field $k$ also shares the characteristic $p$. We will denote the length of an $R$-module by $\ell_{R}(-)$. The well-known Hilbert-Samuel multiplicity is a classical invariant of a local ring that quantifies singularity, where the minimum value of 1 corresponds to a smooth point. In the exploration of singularities, we naturally investigate equimultiple points or the equisingularity of deformations, concepts extensively studied in commutative algebra, algebraic geometry and singularity theory.
In 1983, Monsky [18] introduced a refined version of multiplicity specific to Noetherian local rings of prime characteristic, termed the Hilbert-Kunz multiplicity. It’s worth noting that this notion was also introduced by Kunz [16] in 1969, but Monsky provided a more explicit definition. While this invariant remains largely enigmatic, its study has revealed intriguing connections with the classical multiplicity, albeit often exhibiting more intricate behavior. Unlike the Hilbert-Samuel multiplicity, which is always an integer, the Hilbert-Kunz multiplicity may not be. However, this characteristic enables us to regard it as a more nuanced invariant. Similar to the classical multiplicity, it is capable of discerning regular rings. The study of Hilbert-Kunz multiplicity has engaged numerous researchers, as evidenced by works such as [18], [23], [8], and others.
1.1 Definition.
Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $M$ be a finite $R$-module. The Hilbert-Kunz multiplicity $e_{HK}(I,M)$ of $M$ with respect to $I$ is defined as follows:
(1.1) $e_{HK}(I,M):=\lim_{e\to\infty}\frac{\ell_{R}\left(M/I^{[q]}M\right)}{q^{d}},$
where $q=p^{e}$ and $I^{[q]}$ is the ideal generated by $x^{q}$, $x\in I$. If $M$ is a finite $R$-module $M/I^{[q]}M$ has finite lenght. For simplicity, we can put $e_{HK}(I):=e_{HK}(I,R)$ and $e_{HK}(R):=e_{HK}(\mathfrak{m})$, if it doesn’t cause confusion. Notice that the limit of the right-hand side of (1.1) always exists and is a real number; see, e.g., Monsky [18].
In [17], Kunz observed that the number $\ell_{e}(R):=\ell_{R}(R/\mathfrak{m}^{[q]})/q^{d}$ is a reasonable measure for the singularity of the local ring $R$ and gives a numerical measure of the ’singular-ness’ of $R$. In some sense, it is expected that more larger is the value of $e_{HK}(R)-1$, then more severe is the singularity of $R$.
If $R$ is regular, then $e_{HK}(I,R)=\ell_{R}(R/I)$ for every $\mathfrak{m}$-primary ideal $I$ of $R$ and $e_{HK}(R)=1$. Thus, it is natural to ask: ”If $e_{HK}(R)=1$, is $R$ then regular?”. There is a positive answer to the above question, given by Watanabe and Yoshida [22], when $R$ is unmixed, i. e., given an unmixed local ring $(R,\mathfrak{m})$ of characteristic $p>0$, if $e_{HK}(R)=1$, then $R$ is regular. A somewhat simpler proof can be found in [15]. Watanabe and Yoshida [23] also proved that it holds when $R$ is a two-dimensional Cohen-Macaulay local ring.
The main purpose of this paper is to provide new formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings and consequently produce criteria regarding the regularity of such rings. The study of fiber products and idealization rings has recently received special attention (see for instance [2, 7, 12, 13, 19, 20]). One reason to obtain formulas for the Hilbert-Kunz multiplicity of fiber product rings comes from the fact that the structure of the fiber product ring changes severely, depending on the rings considered (see Corollary 3.6 and Proposition 3.7).
Let us now briefly describe the contents of the paper. Section 2 establishes the main definitions and facts for the rest of the paper. Section 3 is devoted to study the Hilbert-Kunz multiplicity of fiber product rings. Some structural results and examples are also given. Motivated by a result of [21], whose obtain as a consequence a formula for the Hilbert-Kunz multiplicity for idealizations of the form $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$, in the Section 4 we generalize and study the Hilbert-Kunz multiplicity of the idealization ring $R\ltimes M$, where $M$ is a finite $R$-module. Some structural results are provided. In the last section, we discuss two conjectures concerning lower bounds for the Hilbert-Kunz multiplicity. In particular, we provide new classes of $d$-dimensional Noetherian local rings of characteristic $p$ whose satisfy a conjecture posed by Watanabe and Yoshida [24], that asks if $e_{HK}(R)$ is bounded by $e_{HK}(R_{p,d})$, where $(R,\mathfrak{m},k)$ is a $d$-dimensional unmixed local ring of characteristic $p>2$ and dimension $d\geq 2$ and $R_{p,d}$ is the hypersurface $k[[X_{0},X_{1},\ldots,X_{d}]]/(X_{0}^{2}+\cdots+X_{d}^{2})$, with $k=\overline{\mathbb{F}_{p}}$.
2. Setup and Basic facts
We will start by recalling some fundamental properties of the Hilbert-Kunz multiplicity. After that, we will introduce the notion of fiber products and idealization rings to study their Hilbert-Kunz multiplicities. We observe that these two constructions work in general, when the ring characteristic is not necessarily positive.
Recall that if $R$ is a Noetherian local ring and $M$ is a finite $R$-module, then $\dim M$ is defined as $\dim M:=\dim(R/\operatorname{Ann}_{R}(M))$. Also, if $R$ is a local domain, then the rank of $M$ is $\operatorname{rk}_{R}M:=\dim_{Q}(Q\otimes_{R}M)$, where $Q$ is the total quotient ring of $R$.
2.1 Proposition (Fundamental properties (cf. [14, 16, 17, 18, 22])).
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$. Let $I,J$ be $\mathfrak{m}$-primary ideals and let $M$ be a finite $R$-module. Then the following statements hold:
  • (1)
    If $I\subseteq J$, then $e_{HK}(I)\geq e_{HK}(J)$.
  • (2)
    $e_{HK}(I)\geq e_{HK}(R)\geq 1$.
  • (3)
    $e_{HK}(I,M)\geq 0$ and equality holds if and only if $\dim M<\dim R.$
  • (4)
    (Additivity in short exact sequences) If $0\longrightarrow L\longrightarrow M\longrightarrow N\longrightarrow 0$ is a short exact sequence of finite $R$-modules, then
    $e_{HK}(I,M)=e_{HK}(I,L)+e_{HK}(I,N).$
  • (5)
    (Associative Formula)
    $e_{HK}(I,M)=\sum_{\mathfrak{p}\in\operatorname{Assh}(R)}e_{HK}(I,R/\mathfrak{p})\cdot\ell_{R_{\mathfrak{p}}}(M_{\mathfrak{p}}).$
    where $\operatorname{Assh}(R)$ denotes the set of prime ideals $\mathfrak{p}$ of $R$ with $\dim R/\mathfrak{p}=\dim R$. In particular, if $R$ is a local domain and $M$ is a torsion-free $R$module of rank $\operatorname{rk}_{R}M$ , then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(R)$
  • (6)
    If $R$ is a local domain, then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(I,R)$.
  • (7)
    If $R$ is regular, then $e_{HK}(I)=\ell_{R}(R/I)$ and $e_{HK}(R)=1$.
  • (8)
    If $x\in I$ is $R$-regular, then $e_{HK}(I)\leq e_{HK}(I/xR)$.
  • (9)
    (Localization) $e_{HK}(R_{\mathfrak{p}})\leq e_{HK}(R)$ for any prime ideal $\mathfrak{p}$ such that $\dim R/\mathfrak{p}+\operatorname{height}\mathfrak{p}=\dim R$
Fiber product rings
The fiber product structure is studied in various categories. In the case of Noetherian local rings, some fundamental properties can be found in [2].
2.2 Definition.
Let $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ be Noetherian local rings, and let $R\stackrel{{\scriptstyle\pi_{R}}}{{\twoheadrightarrow}}T\stackrel{{\scriptstyle\pi_{S}}}{{\twoheadleftarrow}}S$ be surjective homomorphisms of rings. The fiber product
$R\times_{T}S=\{(r,s)\in R\times S\ \mid\ \pi_{R}(r)=\pi_{S}(s)\},$
is a Noetherian local ring with maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}{\mathfrak{m}_{S}}$, residue field $k$, and it is a subring of the usual direct product $R\times S$ (see [2, Lemma 1.2]). Let $\eta_{R}:R\times_{T}S\twoheadrightarrow R$ and $\eta_{S}:R\times_{T}S\twoheadrightarrow S$ be the natural projections $(r,s)\mapsto r$ and $(r,s)\mapsto s$, respectively. Then $R\times_{T}S$ is represented as a pullback diagram:
2.3 Convention.
In general, when we use the symbols $R$, $S$ or $T$, we assume that $R$, $S$ and $T$ are non-zero Noetherian local rings with the maximal ideal $\mathfrak{m}_{R}$, $\mathfrak{m}_{S}$ and $\mathfrak{m}_{T}$, respectively, and with $k$ being the common residue field. Remark that each (finite) module over $R$, $S$ or $T$ acquires a canonical stucture of (finite) $R\times_{T}S$-module via Diagram (2.5).
We will state three already-known facts about fiber product rings to be used in Section 3.
2.4 Fact ([2]).
There is the exact short sequence of $R\times_{T}S$-modules
2.5 Fact ([2], Lemma 1.5).
The following (in)equalitie holds
$\dim(R\times_{T}S)=\max\{\dim R,\dim S\}\geq\min\{\dim R,\dim S\}\geq\dim T$
2.6 Fact.
Let $k$ be a field and let $I$ and $J$ be ideals of $k[[X_{1},\dots,X_{n}]]$ and $k[[Y_{1},\dots,Y_{m}]]$, respectively. Then
$\frac{k[[X_{1},\dots,X_{n}]]}{I}\times_{k}\frac{k[[Y_{1},\dots,Y_{m}]]}{J}\cong\frac{k[[X_{1},\dots,X_{n},Y_{1},\dots,Y_{m}]]}{(I+J+(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m))}.$
Idealization of a module
Next construction was introduced by Nagata and its nice properties were studied by Anderson and Winders [3].
2.7 Definition.
Let $R$ be a commutative ring and let $M$ be a $R$-module. Then $R\ltimes M:=R\times M$ with coordinate wise addition and multiplication given by $(r_{1},m_{1})(r_{2},m_{2}):=(r_{1}r_{2},r_{1}m_{2}+r_{2}m_{1})$ is a commutative ring with unity (1,0) called the idealization of $M$.
2.8 Remark.
Taking a finite $R$-module $M$ over a $d$-dimensional Noetherian local ring of prime chatecteristic $p>0$ with maximal ideal $\mathfrak{m}$ and residue field $k$, it is well known that the idealization ring $R\ltimes M$ is a $d$-dimensional Noetherian local ring of positive chatecteristic $p$ with maximal ideal $\mathfrak{m}\ltimes M$ and residue field k. Also, we can consider $R\hookrightarrow R\ltimes M$ by sending $r\mapsto(r,0)$. Then $R\ltimes M$ is a local $R$-algebra with the same residue field of R. It is discussed in [3].
We will state two already proved theorems about idealization rings to be used in Section 4.
2.9 Theorem ([3], Theorem 3.1).
Let $R$ be a commutative ring, $I$ an ideal of $R$, M an $R$-module and $N$ a submodule of $M$. Then $I\ltimes N$ is an ideal of $R\ltimes M$ if and only if $IM\subseteq N$. When $I\ltimes N$ is an ideal, $M/N$ is an $R/I$-module and $(R\ltimes M)/(I\ltimes M)\cong(R/I)\ltimes(M/N)$. In particular $(R\ltimes M)/(0\ltimes N)\cong R\ltimes(M/N)$ and therefore $(R\ltimes M)/(0\ltimes M)\cong R$. So the ideals of $R\ltimes M$ containing $0\ltimes M$ are of the form $J\ltimes M$ for some ideal $J$ of $R$.
2.10 Theorem ([3], Theorem 3.6).
Let $R$ be a commutative ring and $M$ an $R$-module. Let $I$ be an ideal of $R$ and $N$ a submodule of $M$. Then $I\ltimes N$ is primary if and only if either
  • a)
    $N=M$ and $I$ is a primary ideal of $R$ or
  • b)
    $N\subsetneq M$, $IM\subseteq N$, and $I$ and $N$ are $\mathfrak{p}$-primary where $\mathfrak{p}=\sqrt{I}$
In either case, $I\ltimes N$ is $\sqrt{I}\ltimes M$-primary.
3. Hilbert-Kunz Multiplicity of Fiber Product Rings
In this section, the rings $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ will be considered Noetherian local rings as in the Definition 2.2. Also, assume that $R$ and $S$ have positive characteristic $p$ and $q$, respectively. In this case, the fiber product ring $R\times_{T}S$ is also a Noetherian local ring of positive characteristic $\max\{p,q\}$. Here, fiber product rings $R\times_{T}S$ will be considered under these conditions. To not be confused, note that in this section the rings $R$, $S$, and $T$ will also be considered as $R\times_{T}S$-modules.
3.1 Proposition.
Let $R\times_{T}S$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an $\mathfrak{m}$-primary ideal of $R\times_{T}S$. Then the following equality holds
$e_{HK}(I)=\left\{\begin{array}[]{llll}e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T);&\dim R=\dim S=\dim T\\ e_{HK}(I,R)+e_{HK}(I,S);&\dim R=\dim S>\dim T\\ e_{HK}(I,R);&\dim R>\dim S>\dim T.\end{array}\right.$
Proof.
Using the short exact sequence of $R\times_{T}S$-modules given in Fact 2.4 and the additivity of Hilbert-Kunz multiplicity in exact short sequences, we get
$e_{HK}(I)=e_{HK}(I,R\oplus S)-e_{HK}(I,T).$
Applying the same argument to the exact sequence
we obtain $e(I,R\oplus S)=e(I,R)+e(I,S)$ and then
$e_{HK}(I)=e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T)$
Note that the dimensions of $R$, $S$ and $T$ as $R\times_{T}S$-modules coincide with their respective Krull dimensions. Finally, by using the formula for $\dim(R\times_{T}S)$ given in Fact 2.5 and item (3) of Proposition 2.1 we get the desired formulas. ∎
Next corollary provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
3.2 Corollary.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). Then the following equality holds
$e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(R)+e_{HK}(S)-e_{HK}(T);&\dim R=\dim S=\dim T\\ e_{HK}(R)+e_{HK}(S);&\dim R=\dim S>\dim T\\ e_{HK}(R);&\dim R>\dim S>\dim T.\end{array}\right.$
Proof.
As an immediate consequence of Proposition 3.1, we obtain the formulas
(3.4) $e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S)-e_{HK}(\mathfrak{m},T);&\dim R=\dim S=\dim T\\ e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S);&\dim R=\dim S>\dim T\\ e_{HK}(\mathfrak{m},R);&\dim R>\dim S>\dim T.\end{array}\right.$
Now, we will discuss why we can replace $e_{HK}(\mathfrak{m},R)$, $e_{HK}(\mathfrak{m},S)$ and $e_{HK}(\mathfrak{m},T)$ in (3.4) by $e_{HK}(R)$, $e_{HK}(S)$ and $e_{HK}(T)$, respectively. Note that $R\times_{T}S$ also has characteristic $p$. In general, if $(A,\mathfrak{m}_{A},k)$ and $(B,\mathfrak{m}_{B},k)$ are $d$-dimensional Noetherian local rings of characteristic $p>0$ and $f:A\to B$ is a surjective ring homomophism, then $f(\mathfrak{m}_{A})=\mathfrak{m}_{B}$ and so it is easy to see that $\mathfrak{m}_{A}^{[q]}B=\mathfrak{m}_{B}^{[q]}B$, with $q=p^{e}$, for each positive integer $e$. Also, as $\ell_{A}(N)=\ell_{B}(N)$, for all $B$-module $N$, we obtain:
$e_{HK}(\mathfrak{m}_{A},B)=\lim_{e\to\infty}\frac{\ell_{A}\left(B/\mathfrak{m}^{[q]}_{A}B\right)}{q^{d}}=\lim_{e\to\infty}\frac{\ell_{B}\left(B/\mathfrak{m}_{B}^{[q]}\right)}{q^{d}}=e_{HK}(B).$
Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas. ∎
3.1 Importance of the formula.
Watanabe and Yoshida asked whether it is possible to determine the infimum of Hilbert-Kunz multiplicities over all unmixed Noetherian local rings of characteristic $p>0$ and dimension $d$ that are not regular. Due to this and the Watanabe-Yoshida conjecture (see Conjecture 5.3), some authors (see, e.g., [24, 6, 4, 1]) have determined lower and upper bounds for some unmixed local rings. However, it is also natural to ask whether it is possible to find lower and upper bounds for the Hilbert-Kunz multiplicity of Noetherian local rings that are not necessarily unmixed. In this sense, by using existing results, Corollary 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring. Below, we provide one example among many to illustrate this fact:
3.3 Example.
Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) with $d:=\dim(R)\geq 2$ and $\dim R\geq\dim S>\dim T$. Suppose that $R$ is a formally unmixed local ring that is not regular. Then
$e_{HK}(R\times_{T}S)\geq 1+\frac{1}{d!(d(d+1)+1)^{d}}.$
The inequality follows easily from Corollary 3.2 and [1, Theorem 4.12].
Amalgamated duplication
Let $R$ be a Noetherian local ring. If we assume $\pi_{R}=\pi_{S}:R\to R/I$ (Diagram (2.5)) are the canonical maps and $I$ is a proper $R$-ideal, there is an isomorphism $R\times_{R/I}R\cong R\bowtie I$, where $R\bowtie I$ is the amalgamated duplication of $R$ along $I$, introduced by D’Anna [9]. Remark that, if $R$ is local with maximal ideal $\mathfrak{m}$, then $R\bowtie I$ is a local ring with maximal ideal denoted by $\mathfrak{m}\bowtie I.$ Some structural properties of $R\bowtie I$ are summarized in [2, Theorem 1.8 and 1.9 - 1.9.4].
We can use Corollary 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
3.4 Corollary.
Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an ideal of $R$. If $\dim R/I<\dim R$, then
$e_{HK}(R\bowtie I)=2e_{HK}(R).$
Proof.
It follows directly by Corollary 3.2. ∎
3.5 Remark.
The formula given in Corollary 3.4 generalizes a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an amalgameted duplication of $R$ along an $\mathfrak{m}$-primary ideal $I$. In this case $\dim R/I=0$ and then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
The study of structural properties of fiber product rings has been addressed by several authors, as [2] and [20]. Next corollary provides some structural properties about fiber product rings.
3.6 Corollary.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) and let $I$ be an ideal of $R$. Then
  • (1)
    If $\dim R=\dim S=\dim T$ and $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)>1$, then $R\times_{T}S$ is not a regular local ring.
  • (2)
    If $\dim R=\dim S>\dim T$, then $R\times_{T}S$ is not a regular local ring. In particular, if $\dim R>\dim R/I$, then $R\bowtie I$ is not a regular local ring.
  • (3)
    If $\dim R>\dim S>\dim T$ and $e_{HK}(R)>1$, then $R\times_{T}S$ is not a regular local ring.
Proof.
It is well known that regular local rings have Hilbert-Kunz multiplicity equal to one. In all the above-mentioned cases, using Corollary 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular. ∎
As we showed above, in general, fiber product rings are not regular. The next result provides an equivalent condition for some fiber product rings to be regular.
3.7 Proposition.
Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). If one of the following conditions holds:
  • (1)
    $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings.
  • (2)
    $R\times_{T}S$ is unmixed.
Then $R\times_{T}S$ is regular if and only if $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
Proof.
  • (1)
    By Corollary 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$. Also, we observe that as $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings, then the same is true for $R\times_{T}S$, by [2, Proposition 1.7]. For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [23, Theorem 2.5]). Then $R\times_{T}S$ is regular if and only if
    $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
  • (2)
    The proof follows by Corollary 3.2 and [22, Theorem 1.5].
3.8 Corollary.
Let $R$, $S$ and $T$ be two-dimensional regular local rings of characteristic $p>0$ as in (2.5). Then $R\times_{T}S$ is regular.
To end this section, we will use the formulas given in Corollary 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
3.9 Example.
Let $k$ be a field of characteristic $p>0$. If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then $R\times_{k}S\cong k[[\underline{X},\underline{Y}]]/(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)$, by Fact 2.6. Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
(3.5) $e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)}\right)=e_{HK}(K[[\underline{X}]])=1,$
by Corollary 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
$e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i,j\leq n)}\right)=e_{HK}(K[[\underline{X}]])+e_{HK}(K[[\underline{Y}]])=2,$
by Corollary 3.2.
3.10 Example.
Let $k$ be a field of characteristic $p\geq 3$. Consider the hypersurfaces and $R=k[[X,Y,Z]]/(XY+Z^{n+1})$ and $S=k[[W,U,V]]/(W^{2}+UV^{2}+U^{n-1})$, where $n\in\mathbb{N}$. Then
$R\times_{k}S\cong\frac{k[[X,Y,Z,W,U,V]]}{(XW,XU,XV,YW,YU,YV,ZW,ZU,ZV,XY+Z^{n+1},W^{2}+UV+U^{n-1})},$
by Fact 2.6. In this case, $\dim R=\dim S>\dim k$ and by using the formulas given in [14, Example 3.18], we have
$e_{HK}(R\times_{k}S)$ $=e_{HK}(R)+e_{HK}(S)=\left(2-\frac{1}{n+1}\right)+\left(2-\frac{1}{4(n-2)}\right)$
$=4-\frac{5n-7}{4(n+1)(n-2)},$
for all positive integer $n\geq 4$, by Corollary 3.2.
3.11 Example.
Let $k$ be a field of characteristic $p>0$. Let $R=k[[X]]$, $S=k[[Y]]$ and $T=K[[Z]]/(Z^{n})$, where $n\in\mathbb{N}$ and $n\geq 2$. Then $R\times_{T}S\cong k[[X,Y]](X^{n}Y-Y^{2})$, by [5, Corollary 4.2.7], taking $A=k[[X,Y]](X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$. In this case, $\dim R=\dim S>\dim T$ and we have
$e_{HK}\left(\frac{k[[X,Y]]}{(X^{n}Y-Y^{2})}\right)=e_{HK}(k[[X]])+e_{HK}(k[[Y]])=2,$
for all positive integer $n\geq 2$, by Corollary 3.2.
4. Hilbert-Kunz multiplicity of idealization rings
We will start by proving the following lemma.
4.1 Lemma.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be an $R$-module. If $I$ is an ideal of $R$, $N$ is a submodule of $M$ and $J=I\ltimes N$ is an ideal of $R\ltimes M$, or equivalently $IM\subseteq N$, then
$J^{[p^{e}]}=I^{[p^{e}]}(R\ltimes M),$
for all positive integer $e$.
Proof.
First of all, we will fix $q:=p^{e}$. Note that, for $(r,m)\in R\ltimes M$ we have that $(r,m)^{n}=(r^{n},nr^{n-1}m)$, for all positive integer $n$. So
$(r,m)^{q}=(r^{q},qr^{q-1}m)=(r^{q},p^{e}r^{q-1}m)=(r^{q},0).$
Then the generators of the ideal $J^{[q]}$ are of the form $(r^{q},0)$, where $r\in I$, and its elements have the form $\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)$ where $r_{i}\in I$, $s_{i}\in R$ and $m_{i}\in M$, for $i=1,\dots,n$. Observe that
$\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)=\sum_{i=1}^{n}(r_{i}^{q},0)(s_{i},m_{i})\in I^{[q]}(R\ltimes M).$
On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$. Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we get
$r(s,m)=\sum_{i=1}^{n}y_{i}x_{i}^{q}(s,m)=\sum_{i=1}^{n}(y_{i}s,y_{i}m)(x_{i}^{q},0)\in J^{[q]}.$
The next theorem provides as a consequence a formula for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module.
4.2 Theorem.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. If one of the following conditions holds:
  • (1)
    $I$ is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$.
  • (2)
    $I$ is an $\mathfrak{m}$-primary ideal of $R$, $N\subsetneq M$ is a submodule of $M$ which is $\mathfrak{m}$-primary and such that $J=I\ltimes N$ is an ideal of $R\ltimes M$ , or equivalently $IM\subseteq N$.
Then following equalities holds
$e_{HK}(J,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M)=\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}\ltimes I^{[q]}M)}{q^{d}}.$
In particular, if $\dim M<\dim R$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,R)$.
Proof.
Note that in both cases the ideal $J$ is $\mathfrak{m}\ltimes M$-primary, by Theorem, 2.10. Fix $q:=p^{e}$. Using the canonical short exact sequence of $R$-modules
we get
$e_{HK}(I,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M).$
Observe that $\ell_{R}(P)=\ell_{R\ltimes M}(P)$, for each $R\ltimes M$-module $P$. Using Lemma 4.1 we have $J^{[q]}=I^{[q]}(R\ltimes M)$ and then
$e_{HK}(J,R\ltimes M)$ $=\lim_{e\to\infty}\frac{\ell_{R\ltimes M}(R\ltimes M/J^{[q]})}{q^{d}}$
$=\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}(R\ltimes M))}{q^{d}}$
$=e_{HK}(I,R\ltimes M)$
$=e_{HK}(I,R)+e_{HK}(I,M).$
Also, by Theorem 2.9, we obtain the isomorphism of $R$-algebras
$\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\cong\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}$
and an exact sequence of $R$-modules
[…]
Using length additivity in short exact sequences we obtain
$\ell_{R}\left(\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\right)=\ell_{R}\left(\frac{R}{I^{[q]}}\right)+\ell_{R}\left(\frac{M}{I^{[q]}M}\right).$
Now, taking limits we get the equality:
$\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}\ltimes I^{[q]}M)}{q^{d}}$ $=\lim_{e\to\infty}\frac{\ell_{R}(R/I^{[q]})}{q^{d}}+\lim_{e\to\infty}\frac{\ell_{R}(M/I^{[q]}M)}{q^{d}}$
$=e_{HK}(I,R)+e_{HK}(I,M).$
4.3 Corollary.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
  • (1)
    $e_{HK}(R\ltimes M)=e_{HK}(R)+e_{HK}(\mathfrak{m},M)$. In particular, if $\dim M<\dim R$, then $e_{HK}(R\ltimes M)=e_{HK}(R)$.
  • (2)
    If $I$ is an ideal of $R$ and $\dim R/I<\dim R$, then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
Proof.
  • (1)
    It is immediate from Theorem 2.9, by taking $J=\mathfrak{m}\ltimes R$.
  • (2)
    Using the canonical short exact sequence of $R$-modules
    we get $e_{HK}(R)=e_{HK}(\mathfrak{m},I)+e_{HK}(\mathfrak{m},R/I)$. Since $\dim R/I<\dim R$, then $e_{HK}(\mathfrak{m},R/I)=0$ and $e_{HK}(R)=e_{HK}(\mathfrak{m},I)$. Utilizing item (1) we obtain the desired equality.
4.4 Remark.
The formulas given in Corollary 4.3 generalize a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$. In this case $\dim R/I=0$ and then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
The next corollary concerns providing formulas for the Hilbert-Kunz multiplicity $e_{HK}(J,R\ltimes M)$, where $J$ is an ideal of $R\ltimes M$ as in Theorem 2.9, by using structural and homological properties of $M$.
4.5 Corollary.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $J$ be an ideal as in Theorem 2.9. Then
  • (1)
    $e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
  • (2)
    If $\operatorname{pd}_{R}M=n<\infty$, then $e_{HK}(J,R\ltimes M)=\left(\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)+1\right)e_{HK}(I,R)$, where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
  • (3)
    If $R$ is a local domain, then $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R)$. Also, if we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
Proof.
  • (1)
    Considering a short exact sequence
    we get
    $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M).$
    Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$. From this inequality and by Theorem 4.2, we obtain the result.
  • (2)
    Taking a minimal free-resolution of $M$
    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences we obtain
    $e_{HK}(I,M)=\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)e_{HK}(I,R).$
    Utilizing Theorem 2.9 we obtain the desired equality.
  • (3)
    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
4.6 Example.
Let $R$ be a Noetherian local ring of dimension $d>0$ and characteristic $p>0$. Then
$e_{HK}\left(\frac{R[X_{1},\dots,X_{r}]}{(X_{1},\dots,X_{r})^{2}}\right)=(r+1)e_{HK}(R).$
Indeed, if $M=R^{r}$, then $R\ltimes M\cong R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2], and the equality follows by the above corollary.
The next corollary provides some important structural information, in the sense of studying the regularity of local rings.
4.7 Corollary.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
  • (1)
    If $\dim M=\dim R$ or $e_{HK}(R)>1$, then $R\ltimes M$ is not a regular local ring.
  • (2)
    If $R$ is unmixed and there exists a finite $R$-module $M$ with $e_{HK}(R\ltimes M)=1$, then $\dim M<\dim R$ and $R$ is a regular local ring.
Proof.
  • (1)
    By Corollary 4.3, we have $e_{HK}(R\ltimes M)=e_{HK}(R)+e_{HK}(\mathfrak{m},M)>1,$ as we suppose $\dim M=\dim R$ or $e_{HK}(R)>1$. Then $R\ltimes M$ can not be regular.
  • (2)
    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [22, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
5. Some bounds for the Hilbert-Kunz multiplicity
There are some fundamental conjectures about Hilbert-Kunz multiplicity to be answered. The following conjecture of Watanabe and Yoshida can be found in [22].
5.1 Conjecture.
Let $R$ be a Cohen-Macaulay local ring of characteristic $p>0$. Then for any $\mathfrak{m}$-primary ideal $I$, we have
  • (1)
    $e_{HK}(I,R)\geq\ell_{R}(R/I)$.
  • (2)
    If $\operatorname{pd}_{R}(R/I)<\infty$, then $e_{HK}(I,R)=\ell_{R}(R/I)$.
It is well known that the above conjecture holds if $R$ is a complete intersection (see Dutta [10, Theorem 1.9]).
The next proposition extends the result already proved for complete intersections to an idealization ring $R\ltimes M$ where $M$ is a finite $R$-module, by using Theorem 4.2. It consists of obtaining an inequality as in the item (1) of Conjecture 5.1 proposed by Watanabe and Yoshida.
5.2 Proposition.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. If $R$ is a complete intersection, $I$ is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$, then
$e_{HK}(J,R\ltimes M)\geq\ell_{R\ltimes M}\left(\frac{R\ltimes M}{J}\right).$
Proof.
Since $R$ is a complete intersection, then it is true that $e_{HK}(I,R)\geq\ell_{R}(R/I)$, by [10, Theorem 1.9]. Using Theorem 4.2, we obtain
$e_{HK}(J,R\ltimes M)\geq e_{HK}(I,R)\geq\ell_{R}(R/I)$
and by Theorem 2.9 we have a canonical isomorphism $R\ltimes M/J\cong R/I$. Therefore, we have
$e_{HK}(J,R\ltimes M)\geq\ell_{R}\left(\frac{R}{I}\right)=\ell_{R\ltimes M}\left(\frac{R\ltimes M}{J}\right).$
New estimates on Watanabe-Yoshida Conjecture
As we observe in Section 1, it is well known that study of Hilbert-Kunz multiplicity is an important tool to investigate the regularity and singularity of unmixed local rings. In this sense, it is natural to ask:
Question: Is there a constant $\epsilon(d)$, depending only on the dimension $d$ of $R$, such that if $R$ is an unmixed local ring which is not regular, then $e_{\text{HK}}(R)-1\geq\epsilon(d)$? Aberbach and Enescu provided a positive answer to this question ([1]). However, if we remove the unmixed condition, does the answer remain valid?
The question above was made even more explicit when Watanabe and Yoshida posed the following conjecture.
5.3 Conjecture (Watanabe-Yoshida [24]).
Let $d\geq 2$ be an integer and $p>2$ be a prime number. Set
$R_{p,d}:=\frac{k[[X_{0},X_{1},\ldots,X_{d}]]}{(X_{0}^{2}+\cdots+X_{d}^{2})}.$
Let $(R,\mathfrak{m},k)$ be a $d$-dimensional unmixed local ring of characteristic $p$ and let $k=\overline{\mathbb{F}_{p}}$. Then the following statements hold:
  1. (1)
    If $R$ is not regular, then $e_{\text{HK}}(R)\geq e_{\text{HK}}(R_{p,d})\geq 1+m_{d}$.
  2. (2)
    If $e_{\text{HK}}(R)=e_{\text{HK}}(R_{p,d})$, then the $\mathfrak{m}$-adic completion of $R$ is isomorphic to $R_{p,d}$ as local rings.
Here, the numbers $m_{d}$ are defined by the relation
$\sec x+\tan x=1+\sum_{d=1}^{\infty}m_{d}x^{d}\quad(-\pi/2<x<\pi/2).$
Recently, Cox and Aberbach [8] proved that the Watanabe-Yoshida conjecture holds up to dimension $7$. Also, the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11]
In our work, we calculated their Hilbert-Kunz multiplicity and gave conditions for fiber products and idealization rings not to be regular. The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas maded in the sections 3 and 4.
5.4 Proposition.
Let $d\geq 2$ be an integer and let $p>2$ be a prime number. Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p$ as in (2.5) and assume $R$ is a $d$-dimensional complete intersection, not regular. If $\dim R\geq\dim S>\dim T$, then
$e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p}).$
Proof.
The proof follows by Corollary 3.2 and [11, Theorem 4.6]. ∎
5.5 Corollary.
Let $d\geq 2$ be an integer, $p>2$ be a prime number, $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular, and let $I$ be an ideal of $R$. If $\dim R/I<\dim R$, then
$e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p}).$
5.6 Proposition.
Let $d\geq 2$ be an integer, $p>2$ be a prime number and let $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular. If $M$ is a finite $R$-module, then
$e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p}).$
Proof.
The proof follows by Corollary 4.3 and [11, Theorem 4.6]. ∎
5.7 Remark.
Propositions 5.2 and 5.4, 5.6 effectively expand the class of rings and ideals that satisfy the desired inequalities of conjectures 5.1 and 5.3, respectively, because an idealization ring $R\ltimes M$ and a fiber product ring $R\times_{T}S$ may not even be Cohen-Macaulay when $R$ is a complete intersection local ring. Both situations are exemplified below.
5.8 Example ([3], Example 4.15).
Let $(R,\mathfrak{m})$ be an $d$-dimensional regular local ring with $x_{1},\dots,x_{d}$ a minimal basis for $\mathfrak{m}$. Let $M_{i}=R/(x_{1},\dots,x_{i})$, $0\leq i\leq d$ (for $i=0$, $M_{i}=R$). Then $\dim R\ltimes M_{i}=d$ and $\operatorname{depth}R\ltimes M_{i}=d-i$. So $R\ltimes M_{i}$ is not Cohen-Macaulay, for $0<i\leq d$.
5.9 Example.
Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$. Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2]. It works event if $R$ and $S$ are regular local rings.

### v3

@@
     (Localization) $e_{HK}(R_{\mathfrak{p}})\leq e_{HK}(R)$ for any prime ideal $\mathfrak{p}$ such that $\dim R/\mathfrak{p}+\operatorname{height}\mathfrak{p}=\dim R$
+  • (10)
+    If $(A,\mathfrak{m})\to(B,\mathfrak{n})$ is a flat local ring homomorphism such that $B/\mathfrak{m}B$ is a field, then $e_{HK}(I)=e_{HK}(IB)$.
 Fiber product rings
@@
 In this case, the fiber product ring $R\times_{T}S$ is also a Noetherian local ring of positive characteristic $\max\{p,q\}$.
-Here, fiber product rings $R\times_{T}S$ will be considered under these conditions.
+During this section, fiber product rings $R\times_{T}S$ will be considered under these conditions.
 To not be confused, note that in this section the rings $R$, $S$, and $T$ will also be considered as $R\times_{T}S$-modules.
-3.1 Proposition.
+3.1 Lemma.
 Let $R\times_{T}S$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an $\mathfrak{m}$-primary ideal of $R\times_{T}S$. Then the following equality holds
@@
 ∎
-Next corollary provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
-3.2 Corollary.
+Next proposition provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
+3.2 Proposition.
 Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). Then the following equality holds
@@
 Proof.
-As an immediate consequence of Proposition 3.1, we obtain the formulas
+As an immediate consequence of Lemma 3.1, we obtain the formulas
 (3.4) $e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S)-e_{HK}(\mathfrak{m},T);&\dim R=\dim S=\dim T\\ e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S);&\dim R=\dim S>\dim T\\ e_{HK}(\mathfrak{m},R);&\dim R>\dim S>\dim T.\end{array}\right.$
@@
 Also, as $\ell_{A}(N)=\ell_{B}(N)$, for all $B$-module $N$, we obtain:
-$e_{HK}(\mathfrak{m}_{A},B)=\lim_{e\to\infty}\frac{\ell_{A}\left(B/\mathfrak{m}^{[q]}_{A}B\right)}{q^{d}}=\lim_{e\to\infty}\frac{\ell_{B}\left(B/\mathfrak{m}_{B}^{[q]}\right)}{q^{d}}=e_{HK}(B).$
-Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas. ∎
+$e_{HK}(\mathfrak{m}_{A},B)$ $=\lim_{e\to\infty}\frac{\ell_{A}\left(B/\mathfrak{m}^{[q]}_{A}B\right)}{q^{d}}$
+$=\lim_{e\to\infty}\frac{\ell_{B}\left(B/\mathfrak{m}_{B}^{[q]}\right)}{q^{d}}$
+$=e_{HK}(B).$
+Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas, by observing that we need only to consider the components of the fiber product $R\times_{T}S$ whose have dimension equal to $\dim R\times_{T}S$.
+∎
 3.1 Importance of the formula.
@@
 However, it is also natural to ask whether it is possible to find lower and upper bounds for the Hilbert-Kunz multiplicity of Noetherian local rings that are not necessarily unmixed.
-In this sense, by using existing results, Corollary 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring.
+In this sense, by using existing results, Proposition 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring.
 Below, we provide one example among many to illustrate this fact:
@@
 $e_{HK}(R\times_{T}S)\geq 1+\frac{1}{d!(d(d+1)+1)^{d}}.$
-The inequality follows easily from Corollary 3.2 and [1, Theorem 4.12].
+The inequality follows easily from Proposition 3.2 and [1, Theorem 4.12].
 Amalgamated duplication
@@
 Remark that, if $R$ is local with maximal ideal $\mathfrak{m}$, then $R\bowtie I$ is a local ring with maximal ideal denoted by $\mathfrak{m}\bowtie I.$ Some structural properties of $R\bowtie I$ are summarized in [2, Theorem 1.8 and 1.9 - 1.9.4].
-We can use Corollary 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
+We can use Proposition 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
 3.4 Corollary.
@@
 Proof.
-It follows directly by Corollary 3.2. ∎
+It follows directly by Proposition 3.2. ∎
 3.5 Remark.
@@
 It is well known that regular local rings have Hilbert-Kunz multiplicity equal to one.
-In all the above-mentioned cases, using Corollary 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular.
+In all the above-mentioned cases, using Proposition 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular.
 ∎
@@
   • (1)
-    By Corollary 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$.
+    By Proposition 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$.
 Also, we observe that as $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings, then the same is true for $R\times_{T}S$, by [2, Proposition 1.7].
-For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [23, Theorem 2.5]).
+For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [24, Theorem 2.5]).
 Then $R\times_{T}S$ is regular if and only if
@@
   • (2)
-    The proof follows by Corollary 3.2 and [22, Theorem 1.5].
+    The proof follows by Proposition 3.2 and [23, Theorem 1.5].
 3.8 Corollary.
 Let $R$, $S$ and $T$ be two-dimensional regular local rings of characteristic $p>0$ as in (2.5). Then $R\times_{T}S$ is regular.
-To end this section, we will use the formulas given in Corollary 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
+To end this section, we will use the formulas given in Proposition 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
 3.9 Example.
-Let $k$ be a field of characteristic $p>0$.
-If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then $R\times_{k}S\cong k[[\underline{X},\underline{Y}]]/(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)$, by Fact 2.6.
-Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
+Let $k$ be a field of characteristic $p>0$. If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then
+$R\times_{k}S\cong\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)},$
+by Fact 2.6. Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
 (3.5) $e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)}\right)=e_{HK}(K[[\underline{X}]])=1,$
-by Corollary 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
+by Proposition 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
 $e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i,j\leq n)}\right)=e_{HK}(K[[\underline{X}]])+e_{HK}(K[[\underline{Y}]])=2,$
-by Corollary 3.2.
+by Proposition 3.2.
 3.10 Example.
@@
 $=4-\frac{5n-7}{4(n+1)(n-2)},$
-for all positive integer $n\geq 4$, by Corollary 3.2.
+for all positive integer $n\geq 4$, by Proposition 3.2.
 3.11 Example.
-Let $k$ be a field of characteristic $p>0$.
-Let $R=k[[X]]$, $S=k[[Y]]$ and $T=K[[Z]]/(Z^{n})$, where $n\in\mathbb{N}$ and $n\geq 2$.
-Then $R\times_{T}S\cong k[[X,Y]](X^{n}Y-Y^{2})$, by [5, Corollary 4.2.7], taking $A=k[[X,Y]](X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$.
-In this case, $\dim R=\dim S>\dim T$ and we have
+Let $k$ be a field of characteristic $p>0$. Let $R=k[[X]]$, $S=k[[Y]]$ and $T=K[[Z]]/(Z^{n})$, where $n\in\mathbb{N}$ and $n\geq 2$. Then
+$R\times_{T}S\cong\frac{k[[X,Y]]}{(X^{n}Y-Y^{2})},$
+by [5, Corollary 4.2.7], taking $A=k[[X,Y]]/(X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$. In this case, $\dim R=\dim S>\dim T$ and we have
 $e_{HK}\left(\frac{k[[X,Y]]}{(X^{n}Y-Y^{2})}\right)=e_{HK}(k[[X]])+e_{HK}(k[[Y]])=2,$
-for all positive integer $n\geq 2$, by Corollary 3.2.
+for all positive integer $n\geq 2$, by Proposition 3.2.
 4. Hilbert-Kunz multiplicity of idealization rings
@@
 $\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)=\sum_{i=1}^{n}(r_{i}^{q},0)(s_{i},m_{i})\in I^{[q]}(R\ltimes M).$
+Then $J^{[q]}\subseteq I^{[q]}(R\ltimes M)$.
 On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$.
@@
 $r(s,m)=\sum_{i=1}^{n}y_{i}x_{i}^{q}(s,m)=\sum_{i=1}^{n}(y_{i}s,y_{i}m)(x_{i}^{q},0)\in J^{[q]}.$
+Then $I^{[q]}(R\ltimes M)\subseteq J^{[q]}$ and the equality holds. ∎
 The next theorem provides as a consequence a formula for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module.
@@
 In this case $\dim R/I=0$ and then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
-The next corollary concerns providing formulas for the Hilbert-Kunz multiplicity $e_{HK}(J,R\ltimes M)$, where $J$ is an ideal of $R\ltimes M$ as in Theorem 2.9, by using structural and homological properties of $M$.
-4.5 Corollary.
-Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module.
-Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $J$ be an ideal as in Theorem 2.9.
-Then
-  • (1)
-    $e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
-  • (2)
-    If $\operatorname{pd}_{R}M=n<\infty$, then $e_{HK}(J,R\ltimes M)=\left(\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)+1\right)e_{HK}(I,R)$, where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$.
-In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
-  • (3)
-    If $R$ is a local domain, then $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R)$.
-Also, if we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
-Proof.
-  • (1)
-    Considering a short exact sequence
-    we get
-    $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M).$
-    Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$. From this inequality and by Theorem 4.2, we obtain the result.
-  • (2)
-    Taking a minimal free-resolution of $M$
-    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences we obtain
-    $e_{HK}(I,M)=\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)e_{HK}(I,R).$
-    Utilizing Theorem 2.9 we obtain the desired equality.
-  • (3)
-    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
-4.6 Example.
-Let $R$ be a Noetherian local ring of dimension $d>0$ and characteristic $p>0$. Then
-$e_{HK}\left(\frac{R[X_{1},\dots,X_{r}]}{(X_{1},\dots,X_{r})^{2}}\right)=(r+1)e_{HK}(R).$
-Indeed, if $M=R^{r}$, then $R\ltimes M\cong R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2], and the equality follows by the above corollary.
-The next corollary provides some important structural information, in the sense of studying the regularity of local rings.
 4.5 Corollary.
@@
   • (2)
-    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [22, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
+    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [23, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
+The next corollary concerns providing formulas for the Hilbert-Kunz multiplicity $e_{HK}(J,R\ltimes M)$, where $J$ is an ideal of $R\ltimes M$ as in Theorem 2.9, by using structural and homological properties of $M$.
+4.6 Corollary.
+Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module.
+Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $J$ be an ideal as in Theorem 2.9.
+Then
+  • (1)
+    $e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
+  • (2)
+    If $\operatorname{pd}_{R}M=n<\infty$, then
+    $e_{HK}(J,R\ltimes M)=\left(\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)+1\right)e_{HK}(I,R),$
+    where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
+  • (3)
+    If $R$ is a local domain, then
+    $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R),$
+    Also, if in addition we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
+Proof.
+  • (1)
+    Considering a short exact sequence
+    we get
+    $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M).$
+    Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$. From this inequality and by Theorem 4.2, we obtain the result.
+  • (2)
+    Taking a minimal free-resolution of $M$
+    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences of finite $R$-modules, we obtain
+    $e_{HK}(I,M)=\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)e_{HK}(I,R).$
+    Utilizing Theorem 2.9 we obtain the desired equality.
+  • (3)
+    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
+4.7 Corollary.
+Let $R$ be a Noetherian local ring of dimension $d>0$ and characteristic $p>0$. Then
+$e_{HK}\left(\frac{R[X_{1},\dots,X_{r}]}{(X_{1},\dots,X_{r})^{2}}\right)=(r+1)e_{HK}(R).$
+Proof.
+If $M=R^{r}$, then $R\ltimes M$ is naturally isomorphic to $R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2]. Therefore, the equality follows by the above corollary. ∎
 5. Some bounds for the Hilbert-Kunz multiplicity
@@
 Recently, Cox and Aberbach [8] proved that the Watanabe-Yoshida conjecture holds up to dimension $7$.
-Also, the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11]
+The first inequality of part (1) of the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11].
+Also, recently, Trivedi [22] proved the second inequality of part (1) of the conjecture, for $d\geq 4$ and $p>d-3$.
 In our work, we calculated their Hilbert-Kunz multiplicity and gave conditions for fiber products and idealization rings not to be regular.
-The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas maded in the sections 3 and 4.
+The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas presented in the sections 3 and 4.
 5.4 Proposition.
@@
 $e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p}).$
+If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the fiber product $R\times_{T}S$.
 Proof.
-The proof follows by Corollary 3.2 and [11, Theorem 4.6]. ∎
+The proof follows by Proposition 3.2, [11, Theorem 4.6] and [22, Theorem (A)]. ∎
 5.5 Corollary.
@@
 $e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p}).$
+If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the amalgameted duplication $R\bowtie I$.
 5.6 Proposition.
@@
 $e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p}).$
+If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the idealization ring $R\ltimes M$.
 Proof.
-The proof follows by Corollary 4.3 and [11, Theorem 4.6]. ∎
+The proof follows by item (1) of Corollary 4.3, [11, Theorem 4.6] and [22, Theorem (A)]
 5.7 Remark.
@@
 5.9 Example.
-Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$.
-Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2].
-It works event if $R$ and $S$ are regular local rings.
+Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$. Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2].
+Competing interests: The authors declare none.

### v4

@@
-Hilbert-Kunz multiplicity of fiber products and idealization rings
+Hilbert-Kunz Multiplicity of Fiber Product Rings and Nagata Idealizations
 Abstract.
-The main purpose of this paper is to provide formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings.
-We calculate the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$, with $R$, $S$ and $T$ being Noetherian local rings with common characteristic and residue field.
-Inspired by a work of Strazzanti and Armengou we calculate the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module.
-We explore certain examples and structural consequences and obtain some bounds for the Hilbert-Kunz multiplicity.
+The main purpose of this paper is to provide formulas for the Hilbert-Kunz multiplicity of fiber product rings and Nagata idealizations.
+We give explicit formulas for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$, where $R$, $S$, and $T$ are Noetherian local rings sharing the same characteristic and residue field.
+We compute the Hilbert-Kunz multiplicity of a general Nagata idealization $R\ltimes M$, where $M$ is a finitely generated $R$-module.
+Additionally, we provide examples, structural results and establish new bounds for the Hilbert-Kunz multiplicity.
 Key words and phrases:
-Hilbert-Kunz, multiplicity, fiber product, idealization
+Fiber product ring, Nagata idealization, Hilbert-Kunz multiplicity, Amalgamated duplication
 2020 Mathematics Subject Classification:
-13D40, 13H15, 13H05
-Corresponding author: P. D. C. Martins
-The first author was supported by grant 2019/21181-0, São Paulo Research Foundation (FAPESP). The second author was supported by grant 2022/12114-0, São Paulo Research Foundation (FAPESP)
+13H15, 13H10
+Corresponding author: Paulo Martins
 1. Introduction
@@
 Similar to the classical multiplicity, it is capable of discerning regular rings.
-The study of Hilbert-Kunz multiplicity has engaged numerous researchers, as evidenced by works such as [18], [24], [8], and others.
 1.1 Definition.
-Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $M$ be a finite $R$-module. The Hilbert-Kunz multiplicity $e_{HK}(I,M)$ of $M$ with respect to $I$ is defined as follows:
+Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $M$ be a finitely generated $R$-module. The Hilbert-Kunz multiplicity $e_{HK}(I,M)$ of $M$ with respect to $I$ is defined as follows:
 (1.1) $e_{HK}(I,M):=\lim_{e\to\infty}\frac{\ell_{R}\left(M/I^{[q]}M\right)}{q^{d}},$
 where $q=p^{e}$ and $I^{[q]}$ is the ideal generated by $x^{q}$, $x\in I$.
-If $M$ is a finite $R$-module $M/I^{[q]}M$ has finite lenght.
+If $M$ is a finitely generated $R$-module $M/I^{[q]}M$ has finite lenght.
 For simplicity, we can put $e_{HK}(I):=e_{HK}(I,R)$ and $e_{HK}(R):=e_{HK}(\mathfrak{m})$, if it doesn’t cause confusion.
 Notice that the limit of the right-hand side of (1.1) always exists and is a real number; see, e.g., Monsky [23].
-In [17], Kunz observed that the number $\ell_{e}(R):=\ell_{R}(R/\mathfrak{m}^{[q]})/q^{d}$ is a reasonable measure for the singularity of the local ring $R$ and gives a numerical measure of the ’singular-ness’ of $R$.
-In some sense, it is expected that more larger is the value of $e_{HK}(R)-1$, then more severe is the singularity of $R$.
-If $R$ is regular, then $e_{HK}(I,R)=\ell_{R}(R/I)$ for every $\mathfrak{m}$-primary ideal $I$ of $R$ and $e_{HK}(R)=1$.
-Thus, it is natural to ask: ”If $e_{HK}(R)=1$, is $R$ then regular?”.
-There is a positive answer to the above question, given by Watanabe and Yoshida [23], when $R$ is unmixed, i.
-e., given an unmixed local ring $(R,\mathfrak{m})$ of characteristic $p>0$, if $e_{HK}(R)=1$, then $R$ is regular.
-A somewhat simpler proof can be found in [15].
-Watanabe and Yoshida [24] also proved that it holds when $R$ is a two-dimensional Cohen-Macaulay local ring.
-The main purpose of this paper is to provide new formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings and consequently produce criteria regarding the regularity of such rings.
-The study of fiber products and idealization rings has recently received special attention (see for instance [2, 7, 12, 13, 19, 20]).
-One reason to obtain formulas for the Hilbert-Kunz multiplicity of fiber product rings comes from the fact that the structure of the fiber product ring changes severely, depending on the rings considered (see Corollary 3.6 and Proposition 3.7).
-Let us now briefly describe the contents of the paper.
-Section 2 establishes the main definitions and facts for the rest of the paper.
-Section 3 is devoted to study the Hilbert-Kunz multiplicity of fiber product rings.
-Some structural results and examples are also given.
-Motivated by a result of [21], whose obtain as a consequence a formula for the Hilbert-Kunz multiplicity for idealizations of the form $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$, in the Section 4 we generalize and study the Hilbert-Kunz multiplicity of the idealization ring $R\ltimes M$, where $M$ is a finite $R$-module.
-Some structural results are provided.
-In the last section, we discuss two conjectures concerning lower bounds for the Hilbert-Kunz multiplicity.
-In particular, we provide new classes of $d$-dimensional Noetherian local rings of characteristic $p$ whose satisfy a conjecture posed by Watanabe and Yoshida [25], that asks if $e_{HK}(R)$ is bounded by $e_{HK}(R_{p,d})$, where $(R,\mathfrak{m},k)$ is a $d$-dimensional unmixed local ring of characteristic $p>2$ and dimension $d\geq 2$ and $R_{p,d}$ is the hypersurface $k[[X_{0},X_{1},\ldots,X_{d}]]/(X_{0}^{2}+\cdots+X_{d}^{2})$, with $k=\overline{\mathbb{F}_{p}}$.
+In [22], Kunz observed that the number $\ell_{e}(R)=\ell_{R}(R/\mathfrak{m}^{[q]})/q^{d}$ is a reasonable measure for the singularity of the local ring $R$ and gives a numerical measure of the ’singular-ness’ of $R$.
+If $R$ is a regular local ring, then $e_{HK}(R)=1$.
+Otherwise, it is expected that more larger is the value of $e_{HK}(R)-1$, then more severe is the singularity of $R$.
+The study of fiber products and idealization rings has recently garnered significant attention (see, for instance, [4, 3, 15, 16, 17, 18, 24, 26]).
+Freitas and Perez [18] obtained explicit formulas for the Hilbert multiplicity of graded fiber product rings and Strazzanti and Armengou [27] obtained, as a consequence of their results for some quadratic quotients of the Rees algebra, a formula for the Hilbert-Kunz multiplicity of Nagata idealizations of the form $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$.
+These results motivated us to obtain explicit formulas for the Hilbert-Kunz multiplicity of a fiber product ring $R\times_{T}S$ (Theorem 3.2) and for a general Nagata idealization $R\ltimes M$ (Corollary 4.3).
+These formulas have yielded new results on the structure of fiber product rings, as well as sharper bounds for the Hilbert-Kunz multiplicity of fiber product rings and Nagata idealizations.
+Let us briefly describe the contents of the paper.
+Section 2 establishes our terminology and some preparatory facts that will be needed throughout the paper.
+Motivated by the formulas of Freitas and Perez [18] for the Hilbert multiplicity of graded fiber product rings, in Section 3, we obtain explicit formulas for the Hilbert-Kunz multiplicity of a fiber product ring.
+We provide a simpler proof of a result established by Freitas and Lima [15] that a non-trivial general fiber product ring $R\times_{T}S$ is not regular.
+Additionally, we present a formula for the Hilbert-Kunz multiplicity of the amalgamated duplication $R\bowtie I$ and examples.
+In Section 4, the main focus is to obtain an explicit formula for the Hilbert-Kunz multiplicity of the general Nagata idealization $R\ltimes M$, where $M$ is a finitely generated $R$-module.
+Section 5 is devoted to studying the behavior of singularities of fiber product rings using a Hilbert-Kunz multiplicity approach.
+In the final section, we establish bounds for the Hilbert-Kunz multiplicity of fiber product rings and Nagata idealizations.
+Furthermore, we extend the formula for the Hilbert-Kunz multiplicity of fiber product rings to the more general case of multi-factor fiber products.
+As an application of our results, we identify new classes of singular rings that satisfy the celebrated Watanabe-Yoshida conjecture.
 2. Setup and Basic Facts
-We will start by recalling some fundamental properties of the Hilbert-Kunz multiplicity.
-After that, we will introduce the notion of fiber products and idealization rings to study their Hilbert-Kunz multiplicities.
-We observe that these two constructions work in general, when the ring characteristic is not necessarily positive.
-Recall that if $R$ is a Noetherian local ring and $M$ is a finite $R$-module, then $\dim M$ is defined as $\dim M:=\dim(R/\operatorname{Ann}_{R}(M))$.
-Also, if $R$ is a local domain, then the rank of $M$ is $\operatorname{rk}_{R}M:=\dim_{Q}(Q\otimes_{R}M)$, where $Q$ is the total quotient ring of $R$.
+We begin by recording some fundamental properties of Hilbert-Kunz multiplicity.
+We then introduce the definitions of fiber products and idealization rings, along with relevant facts that will be used to explore their Hilbert-Kunz multiplicities in the next sections.
 2.1 Proposition (Fundamental properties (cf. [20, 21, 22, 23, 29])).
-Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$. Let $I,J$ be $\mathfrak{m}$-primary ideals and let $M$ be a finite $R$-module. Then the following statements hold:
+Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$.
+Let $I,J$ be $\mathfrak{m}$-primary ideals and let $M$ be a finitely generated $R$-module.
+Then the following statements hold:
   • (1)
@@
   • (4)
-    (Additivity in short exact sequences) If $0\longrightarrow L\longrightarrow M\longrightarrow N\longrightarrow 0$ is a short exact sequence of finite $R$-modules, then
+    (Additivity in short exact sequences) If $0\rightarrow L\rightarrow M\rightarrow N\rightarrow 0$ is a short exact sequence of finitely generated $R$-modules, then
     $e_{HK}(I,M)=e_{HK}(I,L)+e_{HK}(I,N).$
@@
     where $\operatorname{Assh}(R)$ denotes the set of prime ideals $\mathfrak{p}$ of $R$ with $\dim R/\mathfrak{p}=\dim R$.
-In particular, if $R$ is a local domain and $M$ is a torsion-free $R$module of rank $\operatorname{rk}_{R}M$ , then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(R)$
+In particular, if $R$ is a local domain and $M$ is a torsion-free $R$-module of rank $\operatorname{rk}_{R}M$ , then $e_{HK}(I,M)=\operatorname{rk}_{R}M\cdot e_{HK}(I,R)$ .
   • (6)
@@
   • (8)
-    If $x\in I$ is $R$-regular, then $e_{HK}(I)\leq e_{HK}(I/xR)$.
-  • (9)
-    (Localization) $e_{HK}(R_{\mathfrak{p}})\leq e_{HK}(R)$ for any prime ideal $\mathfrak{p}$ such that $\dim R/\mathfrak{p}+\operatorname{height}\mathfrak{p}=\dim R$
-  • (10)
-    If $(A,\mathfrak{m})\to(B,\mathfrak{n})$ is a flat local ring homomorphism such that $B/\mathfrak{m}B$ is a field, then $e_{HK}(I)=e_{HK}(IB)$.
-Fiber product rings
+    If $f:(A,\mathfrak{m})\to(B,\mathfrak{n})$ is a flat local ring homomorphism such that $B/\mathfrak{m}B$ is a field, then $e_{HK}(I)=e_{HK}(IB)$.
+The study of lower bounds for the Hilbert-Kunz multiplicity has been explored by several authors (see e.g., [29, 30, 31, 8, 1]).
+Using an elegant argument, Aberbach and Enescu [1] resolved the fundamental problem of obtaining bounds that are independent of the characteristic in all dimensions.
+We say that a local ring $(A,\mathfrak{m}$) is unmixed (resp.
+quasi-unmixed) if $\dim\widehat{A}/\mathfrak{p}=\dim\widehat{A}$ for every associated (resp.
+minimal) prime ideal of $\widehat{A}$.
+Here, $\widehat{A}$ denotes the $\mathfrak{m}$-adic completion of $A$.
+2.2 Theorem (Aberbach-Enescu).
+Let $(R,\mathfrak{m},k)$ be an unmixed ring of characteristic $p$ and dimension $d\geq 2$. If $R$ is not regular, then
+$e_{HK}(R)\geq 1+\frac{1}{d(d!(d-1)+1)^{d}}.$
+2.1. Fiber product rings
 The fiber product structure is studied in various categories. In the case of Noetherian local rings, some fundamental properties can be found in [4].
@@
 $R\times_{T}S=\{(r,s)\in R\times S\ \mid\ \pi_{R}(r)=\pi_{S}(s)\},$
-is a Noetherian local ring with maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}{\mathfrak{m}_{S}}$, residue field $k$, and it is a subring of the usual direct product $R\times S$ (see [2, Lemma 1.2]).
+is a Noetherian local ring with maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}{\mathfrak{m}_{S}}$, residue field $k$, and it is a subring of the usual direct product $R\times S$ (see [4, Lemma 1.2]).
 Let $\eta_{R}:R\times_{T}S\twoheadrightarrow R$ and $\eta_{S}:R\times_{T}S\twoheadrightarrow S$ be the natural projections $(r,s)\mapsto r$ and $(r,s)\mapsto s$, respectively.
@@
 In general, when we use the symbols $R$, $S$ or $T$, we assume that $R$, $S$ and $T$ are non-zero Noetherian local rings with the maximal ideal $\mathfrak{m}_{R}$, $\mathfrak{m}_{S}$ and $\mathfrak{m}_{T}$, respectively, and with $k$ being the common residue field.
-Remark that each (finite) module over $R$, $S$ or $T$ acquires a canonical stucture of (finite) $R\times_{T}S$-module via Diagram (2.5).
-We will state three already-known facts about fiber product rings to be used in Section 3.
+The fiber product ring $R\times_{T}S$ is non-trivial, i.e., $R\neq T\neq S$, and their maximal ideal is denoted by $\mathfrak{m}:=\mathfrak{m}_{R}\times_{\mathfrak{m}_{T}}\mathfrak{m}_{S}$.
+Note that each (finitely generated) module over $R$, $S$ or $T$ acquires a canonical structure of (finitely generated) $R\times_{T}S$-module via Diagram (2.5).
+We will state known facts about fiber product rings to be used in the next sections.
 2.5 Fact ([4]).
-There is the exact short sequence of $R\times_{T}S$-modules
+There is the short exact sequence of $R\times_{T}S$-modules
+$0\rightarrow R\times_{T}S\rightarrow R\oplus S\rightarrow T\rightarrow 0$
 2.6 Fact ([4], Lemma 1.5).
@@
 $\frac{k[[x_{1},\dots,x_{n}]]}{I}\times_{k}\frac{k[[y_{1},\dots,y_{m}]]}{J}\cong\frac{k[[x_{1},\dots,x_{n},y_{1},\dots,y_{m}]]}{(I+J+(x_{i}y_{j}:1\leq i\leq n,1\leq j\leq m))}.$
-Idealization of a module
-Next construction was introduced by Nagata and its nice properties were studied by Anderson and Winders [3].
+2.8 Fact.
+([26, Fact 2.9]) Let $P=R\times_{k}S$ be a non-trivial fiber product. Then
+$\operatorname{Spec}P=\{\mathfrak{m}_{R}\oplus\mathfrak{q}\mid\mathfrak{q}\in\operatorname{Spec}S\}\cup\{\mathfrak{p}\oplus\mathfrak{m}_{S}\mid\mathfrak{p}\in\operatorname{Spec}R\}.$
+Following this notation, we have $\operatorname{ht}(\mathfrak{m}_{R}\oplus\mathfrak{q})=\operatorname{ht}(\mathfrak{q})$ and $P/(\mathfrak{m}_{R}\oplus\mathfrak{q})\cong S/\mathfrak{q}$, so we must then have $\dim(P/(\mathfrak{m}_{R}\oplus\mathfrak{q}))=\dim(S/\mathfrak{q})$ and similarly for the prime $\mathfrak{p}\oplus\mathfrak{m}_{S}$.
+Futhermore, if $\mathfrak{q}\neq\mathfrak{m}_{S}$, then $P_{\mathfrak{m}_{R}\oplus\mathfrak{q}}\cong S_{\mathfrak{q}}$, and similarly for $\mathfrak{p}\oplus\mathfrak{m}_{S}$.
+2.2. Idealization of a module
+Next construction was introduced by Nagata, and its nice properties were studied by Anderson and Winders [6].
 2.9 Definition.
@@
 2.10 Remark.
-Taking a finite $R$-module $M$ over a $d$-dimensional Noetherian local ring of prime chatecteristic $p>0$ with maximal ideal $\mathfrak{m}$ and residue field $k$, it is well known that the idealization ring $R\ltimes M$ is a $d$-dimensional Noetherian local ring of positive chatecteristic $p$ with maximal ideal $\mathfrak{m}\ltimes M$ and residue field k.
-Also, we can consider $R\hookrightarrow R\ltimes M$ by sending $r\mapsto(r,0)$.
-Then $R\ltimes M$ is a local $R$-algebra with the same residue field of R.
-It is discussed in [3].
+Let $M$ be a finitely generated $R$-module over a $d$-dimensional Noetherian local ring $R$ of prime characteristic $p>0$ with maximal ideal $\mathfrak{m}$ and residue field $k$.
+It is well known that the idealization ring $R\ltimes M$ is also a $d$-dimensional Noetherian local ring of characteristic $p$ with maximal ideal $\mathfrak{m}\ltimes M$ and residue field $k$.
+Furthermore, we can embed $R$ into $R\ltimes M$ via $r\mapsto(r,0)$, making $R\ltimes M$ a local $R$-algebra with the same residue field as $R$.
+This is discussed in detail in [6].
 We will state two already proved theorems about idealization rings to be used in Section 4.
@@
 3. Hilbert-Kunz Multiplicity of Fiber Product Rings
-In this section, the rings $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ will be considered Noetherian local rings as in the Definition 2.2.
-Also, assume that $R$ and $S$ have positive characteristic $p$ and $q$, respectively.
-In this case, the fiber product ring $R\times_{T}S$ is also a Noetherian local ring of positive characteristic $\max\{p,q\}$.
-During this section, fiber product rings $R\times_{T}S$ will be considered under these conditions.
-To not be confused, note that in this section the rings $R$, $S$, and $T$ will also be considered as $R\times_{T}S$-modules.
+In this section, the main focus is to present a formula for the Hilbert-Kunz multiplicity of the fiber product ring $R\times_{T}S$, where $R$, $S$, and $T$ are Noetherian local rings with the same characteristic $p>0$ and residue field.
+In this section, we consider rings $R$, $S$ and $T$ as in Definition 2.3 and adopt the Convention 2.4.
+To avoid confusion, we note that in this section, $R$, $S$, and $T$ will also be treated as $R\times_{T}S$-modules.
+Without loss of generality (see Fact 2.6), we may assume that $\dim R\times_{T}S=\dim R\geq\dim S\geq\dim T$.
 3.1 Lemma.
-Let $R\times_{T}S$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an $\mathfrak{m}$-primary ideal of $R\times_{T}S$. Then the following equality holds
-$e_{HK}(I)=\left\{\begin{array}[]{llll}e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T);&\dim R=\dim S=\dim T\\ e_{HK}(I,R)+e_{HK}(I,S);&\dim R=\dim S>\dim T\\ e_{HK}(I,R);&\dim R>\dim S>\dim T.\end{array}\right.$
+Let $R\times_{T}S$ be a fiber product ring of characteristic $p>0$ and let $I$ be an $\mathfrak{m}$-primary ideal of $R\times_{T}S$. Then the following equality holds
+$e_{HK}(I)=\left\{\begin{array}[]{llll}e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T);&\dim R=\dim S=\dim T\\ e_{HK}(I,R)+e_{HK}(I,S);&\dim R=\dim S>\dim T\\ e_{HK}(I,R);&\dim R>\dim S\geq\dim T.\end{array}\right$.
 Proof.
-Using the short exact sequence of $R\times_{T}S$-modules given in Fact 2.4 and the additivity of Hilbert-Kunz multiplicity in exact short sequences, we get
+Using the short exact sequence of $R\times_{T}S$-modules given in Fact 2.5 and the additivity of Hilbert-Kunz multiplicity in short exact sequences, we get
 $e_{HK}(I)=e_{HK}(I,R\oplus S)-e_{HK}(I,T).$
-Applying the same argument to the exact sequence
-we obtain $e(I,R\oplus S)=e(I,R)+e(I,S)$ and then
-$e_{HK}(I)=e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T)$
+Applying the same argument to the exact sequence $0\rightarrow R\rightarrow R\oplus S\rightarrow S\rightarrow 0$ we obtain $e_{HK}(I,R\oplus S)=e_{HK}(I,R)+e_{HK}(I,S)$ and then
+$e_{HK}(I)=e_{HK}(I,R)+e_{HK}(I,S)-e_{HK}(I,T).$
 Note that the dimensions of $R$, $S$ and $T$ as $R\times_{T}S$-modules coincide with their respective Krull dimensions.
-Finally, by using the formula for $\dim(R\times_{T}S)$ given in Fact 2.5 and item (3) of Proposition 2.1 we get the desired formulas.
+Finally, by using the formula for $\dim(R\times_{T}S)$ given in Fact 2.6 and item (3) of Proposition 2.1 we get the desired formulas, nothing that the components of the fiber product that have dimension less than $\dim R\times_{T}S$ have Hilbert-Kunz multiplicity with respect to $I$ equal to zero.
 ∎
-Next proposition provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
-3.2 Proposition.
-Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). Then the following equality holds
-$e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(R)+e_{HK}(S)-e_{HK}(T);&\dim R=\dim S=\dim T\\ e_{HK}(R)+e_{HK}(S);&\dim R=\dim S>\dim T\\ e_{HK}(R);&\dim R>\dim S>\dim T.\end{array}\right.$
+3.2 Theorem.
+Let $R$, $S$ e $T$ be Noetherian local rings of characteristic $p>0$. Then the following equality holds
+$e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(R)+e_{HK}(S)-e_{HK}(T);&\dim R=\dim S=\dim T\\ e_{HK}(R)+e_{HK}(S);&\dim R=\dim S>\dim T\\ e_{HK}(R);&\dim R>\dim S\geq\dim T.\end{array}\right$.
 Proof.
-As an immediate consequence of Lemma 3.1, we obtain the formulas
-(3.4) $e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S)-e_{HK}(\mathfrak{m},T);&\dim R=\dim S=\dim T\\ e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S);&\dim R=\dim S>\dim T\\ e_{HK}(\mathfrak{m},R);&\dim R>\dim S>\dim T.\end{array}\right.$
+As an immediate consequence of Lemma 3.1, taking $I=\mathfrak{m}$, we obtain the following formulas
+(3.4) $e_{HK}(R\times_{T}S)=\left\{\begin{array}[]{llll}e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S)-e_{HK}(\mathfrak{m},T);&\dim R=\dim S=\dim T\\ e_{HK}(\mathfrak{m},R)+e_{HK}(\mathfrak{m},S);&\dim R=\dim S>\dim T\\ e_{HK}(\mathfrak{m},R);&\dim R>\dim S\geq\dim T.\end{array}\right$.
 Now, we will discuss why we can replace $e_{HK}(\mathfrak{m},R)$, $e_{HK}(\mathfrak{m},S)$ and $e_{HK}(\mathfrak{m},T)$ in (3.4) by $e_{HK}(R)$, $e_{HK}(S)$ and $e_{HK}(T)$, respectively.
-Note that $R\times_{T}S$ also has characteristic $p$.
-In general, if $(A,\mathfrak{m}_{A},k)$ and $(B,\mathfrak{m}_{B},k)$ are $d$-dimensional Noetherian local rings of characteristic $p>0$ and $f:A\to B$ is a surjective ring homomophism, then $f(\mathfrak{m}_{A})=\mathfrak{m}_{B}$ and so it is easy to see that $\mathfrak{m}_{A}^{[q]}B=\mathfrak{m}_{B}^{[q]}B$, with $q=p^{e}$, for each positive integer $e$.
-Also, as $\ell_{A}(N)=\ell_{B}(N)$, for all $B$-module $N$, we obtain:
+It is easy to see that $R\times_{T}S$ also has characteristic $p$.
+Note that, in general, if $(A,\mathfrak{m}_{A},k)$ and $(B,\mathfrak{m}_{B},k)$ are $d$-dimensional Noetherian local rings of characteristic $p>0$ and $f:A\to B$ is a surjective ring homomorphism, then $f(\mathfrak{m}_{A})=\mathfrak{m}_{B}$ and so $\mathfrak{m}_{A}^{[q]}B=\mathfrak{m}_{B}^{[q]}B$, with $q=p^{e}$, for each positive integer $e$.
+Also, as $\ell_{A}(N)=\ell_{B}(N)$, for all $B$-module NN of finite lenght, we obtain:
 $e_{HK}(\mathfrak{m}_{A},B)$ $=\lim_{e\to\infty}\frac{\ell_{A}\left(B/\mathfrak{m}^{[q]}_{A}B\right)}{q^{d}}$
@@
 $=e_{HK}(B).$
-Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas, by observing that we need only to consider the components of the fiber product $R\times_{T}S$ whose have dimension equal to $\dim R\times_{T}S$.
+Therefore, using the formula for $\dim R\times_{T}S$ from Fact 2.6 together with the surjective ring homomorphisms in diagram (2.5), we may substitute $e_{HK}(\mathfrak{m},R)$, $e_{HK}(\mathfrak{m},S)$, and $e_{HK}(\mathfrak{m},T)$ in (3.4) with $e_{HK}(R)$, $e_{HK}(S)$, and $e_{HK}(T)$, respectively.
+Here we observe that it suffices to consider only those components of the fiber product $R\times_{T}S$ whose dimensions are equal to $\dim R\times_{T}S$.
 ∎
-3.1 Importance of the formula.
-Watanabe and Yoshida asked whether it is possible to determine the infimum of Hilbert-Kunz multiplicities over all unmixed Noetherian local rings of characteristic $p>0$ and dimension $d$ that are not regular.
-Due to this and the Watanabe-Yoshida conjecture (see Conjecture 5.3), some authors (see, e.g., [25, 6, 4, 1]) have determined lower and upper bounds for some unmixed local rings.
-However, it is also natural to ask whether it is possible to find lower and upper bounds for the Hilbert-Kunz multiplicity of Noetherian local rings that are not necessarily unmixed.
-In this sense, by using existing results, Proposition 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring.
-Below, we provide one example among many to illustrate this fact:
+The study of structural properties of fiber product rings has been addressed by several authors, as [17] and [4].
+In [25, Proposition 2.6], they proved that a non-trivial fiber product of the form $R\times_{k}S$ is not regular.
+Freitas and Lima [15, Theorem 4.7(1)] extended this result to any non-trivial fiber product ring using a Betti numbers approach.
+We give a simpler proof that any non-trivial fiber product $R\times_{T}S$ is not regular.
+We will make use of the singularness of fiber products in the next sections.
+3.3 Theorem.
+A non-trivial fiber product $R\times_{T}S$ is not a domain.
+Proof.
+Since $\pi_{R}$ and $\pi_{S}$ are surjective and not isomorphisms, we have $\ker(\pi_{R})\neq 0$ and $\ker(\pi_{S})\neq 0$.
+Let $r_{0}\in\ker(\pi_{R})\setminus\{0\}$ and $s_{0}\in\ker(\pi_{S})\setminus\{0\}$.
+Then,
+$\pi_{R}(r_{0})=0=\pi_{S}(0)\Rightarrow(r_{0},0)\in R\times_{T}S,$
+$\pi_{R}(0)=0=\pi_{S}(s_{0})\Rightarrow(0,s_{0})\in R\times_{T}S.$
+However, $(r_{0},0)\cdot(0,s_{0})=(r_{0}\cdot 0,0\cdot s_{0})=(0,0),$ and both $(r_{0},0)\neq 0$ and $(0,s_{0})\neq 0$. Hence, $R\times_{T}S$ is not a domain. ∎
+3.4 Corollary.
+Let $R$, $S$, and $T$ be Noetherian local rings. Then the fiber product $R\times_{T}S$ is not regular.
+3.1. Amalgamated duplication
+Let $R$ be a Noetherian local ring.
+If we assume $\pi_{R}=\pi_{S}:R\to R/I$ (Diagram (2.5)) are the canonical maps and $I$ is a proper $R$-ideal, then $R\times_{R/I}R=R\bowtie I$, where $R\bowtie I$ is the amalgamated duplication of $R$ along $I$, introduced by D’Anna [11].
+Remark that, if $R$ is local with maximal ideal $\mathfrak{m}$, then $R\bowtie I$ is a local ring with maximal ideal denoted by $\mathfrak{m}\bowtie I.$ Some structural properties of $R\bowtie I$ are summarized in [4, Theorem 1.8 and 1.9 - 1.9.4].
+We can use Theorem 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
+3.5 Corollary.
+Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be a proper ideal of $R$. Then
+$e_{HK}(R\bowtie I)=\left\{\begin{array}[]{llll}2e_{HK}(R)-e_{HK}(R/I);&\dim R=\dim R/I\\ 2e_{HK}(R);&\dim R>\dim R/I.\end{array}\right$.
+In particular, if $I$ contains a non-zero divisor, then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
+Proof.
+The proof follows directly using the equality $R\times_{R/I}R=R\bowtie I$ and the formulas given in Theorem 3.2. ∎
+3.6 Remark.
+The formula given in Corollary 3.5 generalizes a formula obtained in [27, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an amalgamated duplication of $R$ along an $\mathfrak{m}$-primary ideal $I$.
+In this case $\dim R/I=0$ and then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
+Since $R\bowtie I=R\times_{R/I}R$, the next result is a consequence of Theorem 3.3.
+3.7 Corollary.
+Let $R$ be a Noetherian local ring and let $I$ be a proper ideal of $R$. Then $R\bowtie I$ is not regular.
+To end this section, we will use the formulas given in Theorem 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
 3.8 Example.
-Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) with $d:=\dim(R)\geq 2$ and $\dim R\geq\dim S>\dim T$.
-Suppose that $R$ is a formally unmixed local ring that is not regular.
-Then
-$e_{HK}(R\times_{T}S)\geq 1+\frac{1}{d!(d(d+1)+1)^{d}}.$
-The inequality follows easily from Proposition 3.2 and [1, Theorem 4.12].
-Amalgamated duplication
-Let $R$ be a Noetherian local ring.
-If we assume $\pi_{R}=\pi_{S}:R\to R/I$ (Diagram (2.5)) are the canonical maps and $I$ is a proper $R$-ideal, there is an isomorphism $R\times_{R/I}R\cong R\bowtie I$, where $R\bowtie I$ is the amalgamated duplication of $R$ along $I$, introduced by D’Anna [9].
-Remark that, if $R$ is local with maximal ideal $\mathfrak{m}$, then $R\bowtie I$ is a local ring with maximal ideal denoted by $\mathfrak{m}\bowtie I.$ Some structural properties of $R\bowtie I$ are summarized in [2, Theorem 1.8 and 1.9 - 1.9.4].
-We can use Proposition 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
-3.4 Corollary.
-Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be an ideal of $R$. If $\dim R/I<\dim R$, then
-$e_{HK}(R\bowtie I)=2e_{HK}(R).$
-Proof.
-It follows directly by Proposition 3.2. ∎
-3.5 Remark.
-The formula given in Corollary 3.4 generalizes a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an amalgameted duplication of $R$ along an $\mathfrak{m}$-primary ideal $I$.
-In this case $\dim R/I=0$ and then $e_{HK}(R\bowtie I)=2e_{HK}(R)$.
-The study of structural properties of fiber product rings has been addressed by several authors, as [2] and [20]. Next corollary provides some structural properties about fiber product rings.
-3.6 Corollary.
-Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5) and let $I$ be an ideal of $R$. Then
-  • (1)
-    If $\dim R=\dim S=\dim T$ and $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)>1$, then $R\times_{T}S$ is not a regular local ring.
-  • (2)
-    If $\dim R=\dim S>\dim T$, then $R\times_{T}S$ is not a regular local ring. In particular, if $\dim R>\dim R/I$, then $R\bowtie I$ is not a regular local ring.
-  • (3)
-    If $\dim R>\dim S>\dim T$ and $e_{HK}(R)>1$, then $R\times_{T}S$ is not a regular local ring.
-Proof.
-It is well known that regular local rings have Hilbert-Kunz multiplicity equal to one.
-In all the above-mentioned cases, using Proposition 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular.
-∎
-As we showed above, in general, fiber product rings are not regular. The next result provides an equivalent condition for some fiber product rings to be regular.
-3.7 Proposition.
-Let $R$, $S$ and $T$ be Noetherian local rings of characteristic $p>0$ as in (2.5). If one of the following conditions holds:
-  • (1)
-    $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings.
-  • (2)
-    $R\times_{T}S$ is unmixed.
-Then $R\times_{T}S$ is regular if and only if $e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
-Proof.
-  • (1)
-    By Proposition 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$.
-Also, we observe that as $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings, then the same is true for $R\times_{T}S$, by [2, Proposition 1.7].
-For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [24, Theorem 2.5]).
-Then $R\times_{T}S$ is regular if and only if
-    $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)=1.$
-  • (2)
-    The proof follows by Proposition 3.2 and [23, Theorem 1.5].
-3.8 Corollary.
-Let $R$, $S$ and $T$ be two-dimensional regular local rings of characteristic $p>0$ as in (2.5). Then $R\times_{T}S$ is regular.
-To end this section, we will use the formulas given in Proposition 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
-3.9 Example.
-Let $k$ be a field of characteristic $p>0$. If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then
-$R\times_{k}S\cong\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)},$
-by Fact 2.6. Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
-(3.5) $e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)}\right)=e_{HK}(K[[\underline{X}]])=1,$
-by Proposition 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
-$e_{HK}\left(\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i,j\leq n)}\right)=e_{HK}(K[[\underline{X}]])+e_{HK}(K[[\underline{Y}]])=2,$
-by Proposition 3.2.
+To compute the Hilbert–Kunz multiplicity of the fiber product $R^{(r_{1})}\times_{k}R^{(r_{2})}$, where $R^{(r_{1})}$ and $R^{(r_{2})}$ are Veronese subrings of $R=k[[x_{1},\dots,x_{d}]]$, we use Theorem 3.2 together with the explicit formula from [31, (1.1) pg.
+53].
+Recall that a Veronese subring of order $r$ is defined by $R^{(r)}=\bigoplus_{n\geq 0}R_{nr},$ where $R_{nr}$ denotes the homogeneous component of degree $nr$.
+The Hilbert–Kunz multiplicity of the fiber product is then given by:
+$e_{HK}\big{(}R^{(r_{1})}\times_{k}R^{(r_{2})}\big{)}=\frac{1}{r_{1}}\binom{d+r_{1}-1}{r_{1}-1}+\frac{1}{r_{2}}\binom{d+r_{2}-1}{r_{2}-1}$.
+Geometrically, the fiber product $R^{(r_{1})}\times_{k}R^{(r_{2})}$ can be interpreted as the union of two formal affine cones over $\mathrm{Spec}(k)$, each defined by a Veronese subring of different order.
 3.9 Example.
 Let $k$ be a field of characteristic $p\geq 3$. Consider the hypersurfaces and $R=k[[x,y,z]]/(xy+z^{n+1})$ and $S=k[[w,u,v]]/(w^{2}+uv^{2}+u^{n-1})$, where $n\in\mathbb{N}$. Then
-$R\times_{k}S\cong\frac{k[[X,Y,Z,W,U,V]]}{(XW,XU,XV,YW,YU,YV,ZW,ZU,ZV,XY+Z^{n+1},W^{2}+UV+U^{n-1})},$
-by Fact 2.6. In this case, $\dim R=\dim S>\dim k$ and by using the formulas given in [14, Example 3.18], we have
+$R\times_{k}S\cong\frac{k[[x,y,z,w,u,v]]}{(xw,xu,xv,yw,yu,yv,zw,zu,zv,xy+z^{n+1},w^{2}+uv^{2}+u^{n-1})},$
+by Fact 2.7. In this case, $\dim R=\dim S>\dim k$ and by using the formulas given in [20, Example 3.18], we have
 $e_{HK}(R\times_{k}S)$ $=e_{HK}(R)+e_{HK}(S)=\left(2-\frac{1}{n+1}\right)+\left(2-\frac{1}{4(n-2)}\right)$
 $=4-\frac{5n-7}{4(n+1)(n-2)},$
-for all positive integer $n\geq 4$, by Proposition 3.2.
+for all positive integer $n\geq 4$, by Theorem 3.2.
 3.10 Example.
@@
 $R\times_{T}S\cong\frac{k[[x,y]]}{(x^{n}y-y^{2})},$
-by [5, Corollary 4.2.7], taking $A=k[[X,Y]]/(X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$. In this case, $\dim R=\dim S>\dim T$ and we have
+by [9, Corollary 4.2.7], taking $A=k[[x,y]]/(x^{n}y-y^{2})$, $\mathfrak{p}=(x^{n}-y)$ and $\mathfrak{q}=(y)$. In this case, $\dim R=\dim S>\dim T$ and we have
 $e_{HK}\left(\frac{k[[x,y]]}{(x^{n}y-y^{2})}\right)=e_{HK}(k[[x]])+e_{HK}(k[[y]])=2,$
-for all positive integer $n\geq 2$, by Proposition 3.2.
-4. Hilbert-Kunz multiplicity of idealization rings
-We will start by proving the following lemma.
+for all positive integer $n\geq 2$, by Theorem 3.2.
+4. Hilbert-Kunz multiplicity of Nagata idealizations
+We will start by proving the following lemma, which will play a crucial role in proving our main theorem of this section.
 4.1 Lemma.
 Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be an $R$-module.
-If $I$ is an ideal of $R$, $N$ is a submodule of $M$ and $J=I\ltimes N$ is an ideal of $R\ltimes M$, or equivalently $IM\subseteq N$, then
+If $I$ is an ideal of $R$, NN is a submodule of $M$ and $J=I\ltimes N$ is an ideal of $R\ltimes M$, or equivalently $IM\subseteq N$, then
 $J^{[p^{e}]}=I^{[p^{e}]}(R\ltimes M),$
@@
 First of all, we will fix $q:=p^{e}$. Note that, for $(r,m)\in R\ltimes M$ we have that $(r,m)^{n}=(r^{n},nr^{n-1}m)$, for all positive integer $n$. So
-$(r,m)^{q}=(r^{q},qr^{q-1}m)=(r^{q},p^{e}r^{q-1}m)=(r^{q},0).$
-Then the generators of the ideal $J^{[q]}$ are of the form $(r^{q},0)$, where $r\in I$, and its elements have the form $\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)$ where $r_{i}\in I$, $s_{i}\in R$ and $m_{i}\in M$, for $i=1,\dots,n$.
+$(r,m)^{q}=(r^{q},qr^{q-1}m)=(r^{q},p^{e}r^{q-1}m)=(r^{q},0),$
+since $R$ has characteristic $p$.
+Then, the generators of the ideal $J^{[q]}$ are of the form $(r^{q},0)$, where $r\in I$, and its elements have the form $\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)$ where $r_{i}\in I$, $s_{i}\in R$ and $m_{i}\in M$, for $i=1,\dots,n$.
 Observe that
-$\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)=\sum_{i=1}^{n}(r_{i}^{q},0)(s_{i},m_{i})\in I^{[q]}(R\ltimes M).$
+$\sum_{i=1}^{n}(s_{i},m_{i})(r_{i}^{q},0)=\sum_{i=1}^{n}(r_{i}^{q},0)(s_{i},m_{i})=\sum_{i=1}^{n}r_{i}^{q}(s_{i},m_{i})\in I^{[q]}(R\ltimes M).$
 Then $J^{[q]}\subseteq I^{[q]}(R\ltimes M)$.
 On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$.
-Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we get
+Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we have
 $r(s,m)=\sum_{i=1}^{n}y_{i}x_{i}^{q}(s,m)=\sum_{i=1}^{n}(y_{i}s,y_{i}m)(x_{i}^{q},0)\in J^{[q]}.$
 Then $I^{[q]}(R\ltimes M)\subseteq J^{[q]}$ and the equality holds. ∎
-The next theorem provides as a consequence a formula for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes M$, where $M$ is a finite $R$-module.
+The next theorem provides, as a consequence, a formula for the Hilbert-Kunz multiplicity of a general Nagata idealization $R\ltimes M$, where $M$ is a finitely generated $R$-module.
 4.2 Theorem.
-Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. If one of the following conditions holds:
+Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module. If one of the following conditions holds:
   • (1)
-    $I$ is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$.
+    II is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$.
   • (2)
-    $I$ is an $\mathfrak{m}$-primary ideal of $R$, $N\subsetneq M$ is a submodule of $M$ which is $\mathfrak{m}$-primary and such that $J=I\ltimes N$ is an ideal of $R\ltimes M$ , or equivalently $IM\subseteq N$.
-Then following equalities holds
+    II is an $\mathfrak{m}$-primary ideal of $R$, $N\subsetneq M$ is a submodule of $M$ which is $\mathfrak{m}$-primary and such that $J=I\ltimes N$ is an ideal of $R\ltimes M$ , or equivalently $IM\subseteq N$.
+The following equalities hold
 $e_{HK}(J,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M)=\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}\ltimes I^{[q]}M)}{q^{d}}.$
@@
 Proof.
-Note that in both cases the ideal $J$ is $\mathfrak{m}\ltimes M$-primary, by Theorem, 2.10. Fix $q:=p^{e}$. Using the canonical short exact sequence of $R$-modules
-we get
-$e_{HK}(I,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M).$
-Observe that $\ell_{R}(P)=\ell_{R\ltimes M}(P)$, for each $R\ltimes M$-module $P$. Using Lemma 4.1 we have $J^{[q]}=I^{[q]}(R\ltimes M)$ and then
+Note that in both cases the ideal $J$ is $\mathfrak{m}\ltimes M$-primary, by Theorem 2.12. We will fix $q:=p^{e}$. Using the canonical short exact sequence of $R$-modules
+$0\rightarrow R\rightarrow R\ltimes M\rightarrow M\rightarrow 0$
+we get $e_{HK}(I,R\ltimes M)=e_{HK}(I,R)+e_{HK}(I,M)$.
+Observe that $\ell_{R}(P)=\ell_{R\ltimes M}(P)$, for each $R\ltimes M$-module $P$ of finite lenght.
+Also, using Lemma 4.1 we have the equility $J^{[q]}=I^{[q]}(R\ltimes M)$ and then
 $e_{HK}(J,R\ltimes M)$ $=\lim_{e\to\infty}\frac{\ell_{R\ltimes M}(R\ltimes M/J^{[q]})}{q^{d}}$
@@
 $=e_{HK}(I,R)+e_{HK}(I,M).$
-Also, by Theorem 2.9, we obtain the isomorphism of $R$-algebras
-$\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\cong\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}$
-and an exact sequence of $R$-modules
-[…]
-Using length additivity in short exact sequences we obtain
+Also, by Theorem 2.11, we have the isomorphism
+$\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\cong\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}.$
+Then, considering the exact sequence of $R$-modules of finite length
+$0\rightarrow\frac{R}{I^{[q]}}\rightarrow\frac{R}{I^{[q]}}\ltimes\frac{M}{I^{[q]}M}\rightarrow\frac{M}{I^{[q]}M}\rightarrow 0$
+And using length additivity in short exact sequences, we obtain
 $\ell_{R}\left(\frac{R\ltimes M}{I^{[q]}\ltimes I^{[q]}M}\right)=\ell_{R}\left(\frac{R}{I^{[q]}}\right)+\ell_{R}\left(\frac{M}{I^{[q]}M}\right).$
-Now, taking limits we get the equality:
+Now, taking limits, we get the equality:
 $\lim_{e\to\infty}\frac{\ell_{R}(R\ltimes M/I^{[q]}\ltimes I^{[q]}M)}{q^{d}}$ $=\lim_{e\to\infty}\frac{\ell_{R}(R/I^{[q]})}{q^{d}}+\lim_{e\to\infty}\frac{\ell_{R}(M/I^{[q]}M)}{q^{d}}$
@@
 4.3 Corollary.
-Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
+Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module. Then
   • (1)
@@
   • (2)
-    If $I$ is an ideal of $R$ and $\dim R/I<\dim R$, then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
+    If $I$ is a proper ideal of $R$ and $\dim R/I<\dim R$, then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
 Proof.
   • (1)
-    It is immediate from Theorem 2.9, by taking $J=\mathfrak{m}\ltimes R$.
+    It follows directly from Theorem 4.2, by taking $J=\mathfrak{m}\ltimes M$.
   • (2)
-    Using the canonical short exact sequence of $R$-modules
-    we get $e_{HK}(R)=e_{HK}(\mathfrak{m},I)+e_{HK}(\mathfrak{m},R/I)$.
-Since $\dim R/I<\dim R$, then $e_{HK}(\mathfrak{m},R/I)=0$ and $e_{HK}(R)=e_{HK}(\mathfrak{m},I)$.
-Utilizing item (1) we obtain the desired equality.
+    Using the short exact sequence $0\rightarrow I\rightarrow R\rightarrow R/I\rightarrow 0$, we get $e_{HK}(R)=e_{HK}(\mathfrak{m},I)+e_{HK}(\mathfrak{m},R/I)$.
+Since $\dim R/I<\dim R$, then $e_{HK}(\mathfrak{m},R/I)=0$ (see Proposition 2.1(3)) and therefore $e_{HK}(R)=e_{HK}(\mathfrak{m},I)$.
+Utilizing item (1), we obtain the desired equality.
 4.4 Remark.
-The formulas given in Corollary 4.3 generalize a formula obtained in [21, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$.
+The formulas given in Corollary 4.3 generalize a formula obtained in [27, Theorem 2.3 and Remark 2.4] for the Hilbert-Kunz multiplicity of an idealization ring $R\ltimes I$, where $I$ is an $\mathfrak{m}$-primary ideal of $R$.
 In this case $\dim R/I=0$ and then $e_{HK}(R\ltimes I)=2e_{HK}(R)$.
-4.5 Corollary.
-Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module. Then
-  • (1)
-    If $\dim M=\dim R$ or $e_{HK}(R)>1$, then $R\ltimes M$ is not a regular local ring.
-  • (2)
-    If $R$ is unmixed and there exists a finite $R$-module $M$ with $e_{HK}(R\ltimes M)=1$, then $\dim M<\dim R$ and $R$ is a regular local ring.
-Proof.
-  • (1)
-    By Corollary 4.3, we have $e_{HK}(R\ltimes M)=e_{HK}(R)+e_{HK}(\mathfrak{m},M)>1,$ as we suppose $\dim M=\dim R$ or $e_{HK}(R)>1$. Then $R\ltimes M$ can not be regular.
-  • (2)
-    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [23, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
 The next corollary concerns providing formulas for the Hilbert-Kunz multiplicity $e_{HK}(J,R\ltimes M)$, where $J$ is an ideal of $R\ltimes M$ as in Theorem 4.2, by using structural and homological properties of $M$.
 4.5 Corollary.
-Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module.
+Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module.
 Let $I$ be an $\mathfrak{m}$-primary ideal of $R$ and let $J$ be an ideal as in Theorem 4.2.
@@
   • (1)
-    $e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
+    $e_{HK}(I,R)\leq e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
   • (2)
@@
     $e_{HK}(J,R\ltimes M)=\left(\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)+1\right)e_{HK}(I,R),$
-    where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
+    where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, we should have that $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
   • (3)
     If $R$ is a local domain, then
-    $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R),$
-    Also, if in addition we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
+    $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R).$
 Proof.
   • (1)
-    Considering a short exact sequence
-    we get
-    $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M).$
-    Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$. From this inequality and by Theorem 4.2, we obtain the result.
+    Considering a short exact sequence $0\rightarrow K\rightarrow R^{\mu(M)}\rightarrow M\rightarrow 0$ we get $e_{HK}\left(I,R^{\mu(M)}\right)=\mu(M)e_{HK}(I,R)=e_{HK}(I,K)+e_{HK}(I,M)$.
+Then $\mu(M)e_{HK}(I,R)\geq e_{HK}(I,M)$.
+From this inequality and by Theorem 4.2, we obtain the result.
   • (2)
     Taking a minimal free-resolution of $M$
-    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences of finite $R$-modules, we obtain
+    $0\rightarrow R^{\beta_{n}^{R}(M)}\rightarrow\cdots\rightarrow R^{\beta_{1}^{R}(M)}\rightarrow R^{\beta_{0}^{R}(M)}\rightarrow M\rightarrow 0$
+    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences of finitely generated $R$-modules, we obtain
     $e_{HK}(I,M)=\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)e_{HK}(I,R).$
-    Utilizing Theorem 2.9 we obtain the desired equality.
+    Now, utilizing Theorem 4.2 we obtain the desired equality.
   • (3)
-    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
+    The proof follows by Theorem 4.2 and item (6) of Proposition 2.1.
+4.6 Example.
+Let $R$ be a Noetherian local ring of dimension $d>0$ and characteristic $p>0$. Then
+$e_{HK}\left(\frac{R[x_{1},\dots,x_{n}]}{(x_{1},\dots,x_{n})^{2}}\right)=(n+1)e_{HK}(R).$
+Indeed, if $M=R^{n}$ , then $R\ltimes M$ is naturally isomorphic to $R[x_{1},\dots,x_{n}]/(x_{1},\dots,x_{n})^{2}$, by [6, Proposition 2.2]. Therefore, the equality follows by the above corollary.
+The study of structural properties of idealization rings has been addressed by several authors, as [6], [3], and [12].
+The next theorem shows that an idealization ring $R\ltimes M$ is not regular.
+We will make use of the singularness of idealizations in the next sections.
+4.7 Theorem.
+Let $R$ be a Noetherian local ring and $M$ a finitely generated $R$-module. Then $R\ltimes M$ is not regular.
+Proof.
+It suffices to show that $R\ltimes M$ is not a domain.
+In fact, consider $x=(0,m)$ and $y=(0,m^{\prime}),\text{with }m,m^{\prime}\in M,\ m\neq 0,\ m^{\prime}\neq 0.$ Then $x\cdot y=(0,m)(0,m^{\prime})=(0,0\cdot m^{\prime}+0\cdot m)=(0,0).$ Therefore, $R\ltimes M$ has zero divisors and is not a domain.
 4.8 Corollary.
-Let $R$ be a Noetherian local ring of dimension $d>0$ and characteristic $p>0$. Then
-$e_{HK}\left(\frac{R[X_{1},\dots,X_{r}]}{(X_{1},\dots,X_{r})^{2}}\right)=(r+1)e_{HK}(R).$
+Let $A$ be one of the rings $R\times_{T}S$, $R\bowtie I$, or $R\ltimes M$, each of which is a three-dimensional unmixed local ring of characteristic $p>0$. Then $e_{HK}(A)\geq\frac{4}{3}$.
 Proof.
-If $M=R^{r}$, then $R\ltimes M$ is naturally isomorphic to $R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2]. Therefore, the equality follows by the above corollary. ∎
-5. Some bounds for the Hilbert-Kunz multiplicity
-There are some fundamental conjectures about Hilbert-Kunz multiplicity to be answered. The following conjecture of Watanabe and Yoshida can be found in [23].
-5.1 Conjecture.
-Let $R$ be a Cohen-Macaulay local ring of characteristic $p>0$. Then for any $\mathfrak{m}$-primary ideal $I$, we have
+This follows from Corollary 3.4, Theorem 4.7, and [31, Theorem 3.1 (1)]. ∎
+5. Applications to the structure of singularities
+In this section, we give some applications of the results obtained in the previous sections, focusing on the study of the singularities of the unmixed fiber product ring $P=R\times_{k}S$.
+We show, for example, that if $e_{HK}(P)=\frac{8}{3}$, then $P$ is precisely the gluing of two quadratic hypersurfaces in $\mathbb{A}^{4}_{k}$ along the origin.
+Moreover, we prove that if $e_{HK}(P)\in\left(\frac{8}{3},\frac{17}{6}\right],$ then $P$ is exactly the gluing of two hypersurfaces in $\mathbb{A}^{4}_{k}$ along the origin, at least one of type $x^{2}+y^{2}+z^{2}+w^{c}=0,$ for some integer $c\geq 3$.
+These define isolated singularities of type $A$ (or du Val singularities in dimension two), perturbed by a monomial of order at least $3$.
+We also analyze the tangent structure of $P$ at its closed point.
+That is, we consider the associated graded ring $\operatorname{gr}_{\mathfrak{m}}(P)$, which can be interpreted as the first-order approximation of $P$ near the origin-i.e., its tangent cone.
+5.1 Remark.
+Let $A$ be a Noetherian local ring. Then:
+[…]
+Let $(A,\mathfrak{m},k)$ be any local ring of positive dimension. The associated graded ring $\operatorname{gr}_{\mathfrak{m}}(A)$ of $A$ with respect to $\mathfrak{m}$ is defined as follows:
+$\operatorname{gr}_{\mathfrak{m}}(A):=\bigoplus_{n\geq 0}\mathfrak{m}^{n}/\mathfrak{m}^{n+1}.$
+Then $G=\operatorname{gr}_{\mathfrak{m}}(A)$ is a homogeneous $k$-algebra such that $\mathfrak{M}=G_{+}$ is the unique homogeneous maximal ideal of $G$.
+5.2 Facts.
+([5, Remark 2.2]) Consider $P=R\times_{k}S$ with unique maximal ideal $\mathfrak{m}=\mathfrak{m}_{R}\oplus\mathfrak{m}_{S}$. Then
   • (1)
-    $e_{HK}(I,R)\geq\ell_{R}(R/I)$.
+    $\widehat{P}=\widehat{R}\times_{k}\widehat{S}$.
   • (2)
-    If $\operatorname{pd}_{R}(R/I)<\infty$, then $e_{HK}(I,R)=\ell_{R}(R/I)$.
-It is well known that the above conjecture holds if $R$ is a complete intersection (see Dutta [10, Theorem 1.9]).
-The next proposition extends the result already proved for complete intersections to an idealization ring $R\ltimes M$ where $M$ is a finite $R$-module, by using Theorem 4.2.
-It consists of obtaining an inequality as in the item (1) of Conjecture 5.1 proposed by Watanabe and Yoshida.
+    $\operatorname{gr}_{\mathfrak{m}}(P)\cong\operatorname{gr}_{\mathfrak{m}_{R}}(R)\times_{k}\operatorname{gr}_{\mathfrak{m}_{S}}(S)$ .
+5.3 Lemma.
+Let $P=R\times_{k}S$. If $P$ is quasi-unmixed, then $\dim R=\dim S$.
+Proof.
+[…]
+$\operatorname{Spec}\widehat{P}=\operatorname{Spec}\widehat{R}\times_{k}\widehat{S}=\{\mathfrak{m}_{\widehat{R}}\oplus\mathfrak{q}\mid\mathfrak{q}\in\operatorname{Spec}\widehat{S}\}\cup\{\mathfrak{p}\oplus\mathfrak{m}_{\widehat{S}}\mid\mathfrak{p}\in\operatorname{Spec}\widehat{R}\}$.
+and then $\mathfrak{m}_{\widehat{R}}\oplus\mathfrak{q}_{0}$ and $\mathfrak{p}_{0}\oplus\mathfrak{m}_{\widehat{S}}$ are in $\operatorname{Min}(P)$.
+Also, since $\widehat{P}/(\mathfrak{p}_{0}\oplus\mathfrak{m}_{\widehat{S}})\cong\widehat{R}/\mathfrak{p}_{0}$ and $\widehat{P}/(\mathfrak{m}_{\widehat{R}}\oplus\mathfrak{q}_{0})\cong\widehat{S}/\mathfrak{q}_{0}$ and $P$ is quasi-unmixed, we have:
+$\dim\widehat{R}=\dim\widehat{R}/\mathfrak{p}_{0}=\dim\widehat{S}/\mathfrak{q}_{0}=\dim\widehat{S}$.
 5.4 Proposition.
-Let $(R,\mathfrak{m},k)$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finite $R$-module.
-If $R$ is a complete intersection, $I$ is an $\mathfrak{m}$-primary ideal of $R$ and $J=I\ltimes M$, then
-$e_{HK}(J,R\ltimes M)\geq\ell_{R\ltimes M}\left(\frac{R\ltimes M}{J}\right).$
+Let $P=R\times_{k}S$. Then $P$ is unmixed if and only if $R$ and $S$ are unmixed and $\dim R=\dim S$.
 Proof.
-Since $R$ is a complete intersection, then it is true that $e_{HK}(I,R)\geq\ell_{R}(R/I)$, by [10, Theorem 1.9]. Using Theorem 4.2, we obtain
-$e_{HK}(J,R\ltimes M)\geq e_{HK}(I,R)\geq\ell_{R}(R/I)$
-and by Theorem 2.9 we have a canonical isomorphism $R\ltimes M/J\cong R/I$. Therefore, we have
-$e_{HK}(J,R\ltimes M)\geq\ell_{R}\left(\frac{R}{I}\right)=\ell_{R\ltimes M}\left(\frac{R\ltimes M}{J}\right).$
-New estimates on Watanabe-Yoshida Conjecture
-As we observe in Section 1, it is well known that study of Hilbert-Kunz multiplicity is an important tool to investigate the regularity and singularity of unmixed local rings.
-In this sense, it is natural to ask:
-Question: Is there a constant $\epsilon(d)$, depending only on the dimension $d$ of $R$, such that if $R$ is an unmixed local ring which is not regular, then $e_{\text{HK}}(R)-1\geq\epsilon(d)$?
-Aberbach and Enescu provided a positive answer to this question ([1]).
-However, if we remove the unmixed condition, does the answer remain valid?
-The question above was made even more explicit when Watanabe and Yoshida posed the following conjecture.
+By Remark 5.1, it suffices to show the following:
+  • (a)
+    $P$ is quasi-unmixed if and only if $R$ and $S$ are quasi-unmixed and $\dim R=\dim S$.
+  • (b)
+    Let $\dim R=\dim S$. Then $P$ satisfies $(S_{1})$ if and only if $R$ and $S$ satisfy $(S_{1})$.
+Proof of (a): Suppose that $P$ is quasi-unmixed. By Lemma 5.3 and Fact 2.6, we have $\dim P=\dim R=\dim S$. Recall that, we have:
+$\operatorname{Spec}\widehat{P}=\{\mathfrak{m}_{\widehat{R}}\oplus\mathfrak{q}\mid\mathfrak{q}\in\operatorname{Spec}\widehat{S}\}\cup\{\mathfrak{p}\oplus\mathfrak{m}_{\widehat{S}}\mid\mathfrak{p}\in\operatorname{Spec}\widehat{R}\}$.
+[…]
+(5.1) $\widehat{P}/(\mathfrak{p}\oplus\mathfrak{m}_{\widehat{S}})\cong\widehat{R}/\mathfrak{p},\quad\widehat{P}/(\mathfrak{m}_{\widehat{R}}\oplus\mathfrak{q})\cong\widehat{S}/\mathfrak{q},$
+[…]
+[…]
+Proof of (b): Let $\mathfrak{p}\in\operatorname{Spec}P$, $\mathfrak{p}\neq\mathfrak{m}=\mathfrak{m}_{R}\oplus\mathfrak{m}_{S}$.
+Say $\mathfrak{p}=\mathfrak{m}_{R}\oplus\mathfrak{p}_{S}$, where $\mathfrak{p}_{S}\in\operatorname{Spec}(S)$ is not maximal.
+Then $P_{\mathfrak{p}}\cong S_{\mathfrak{p}_{S}}$, by Fact 2.8.
+We have
+$\operatorname{depth}(P_{\mathfrak{p}})=\operatorname{depth}(S_{\mathfrak{p}_{S}})\quad\text{and}\quad\dim(P_{\mathfrak{p}})=\dim(S_{\mathfrak{p}_{S}}).$
+Therefore, in this case, we have
+[…]
+Similarly, if $\mathfrak{p}=\mathfrak{p}_{R}\oplus\mathfrak{m}_{S}$, where $\mathfrak{p}_{R}\in\operatorname{Spec}(R)$ is not maximal. We get
+[…]
+Now, let $\mathfrak{m}=\mathfrak{m}_{R}\oplus\mathfrak{m}_{S}$ be the maximal ideal of $P$. We know that $\dim P=\dim R=\dim S$ (see Fact 2.6) and
+$0pt(P)=\min\{0ptR,0ptS,1\},$
+by [15, Corollary 4.5]. Thus, it is easy to see that $P$ satisfies $(S_{1})$ at $\mathfrak{m}$ if and only if $R$ and $S$ satisfy $(S_{1})$ at $\mathfrak{m}_{R}$ and $\mathfrak{m}_{S}$, respectively.
+5.5 Question.
+Let $P=R\times_{T}S$, where $T\neq k$. Is it true that
+[…]
+5.6 Theorem.
+Let $P=R\times_{k}S$ be a three-dimensional unmixed local ring of characteristic $p>0$, where $R$ and $S$ are non-regular rings of characteristic $p$. Then the following statements hold:
+  1. (1)
+    $e_{HK}(P)\geq\dfrac{8}{3}$.
+  2. (2)
+    Suppose that $k=\overline{k}$ and $\operatorname{char}k\neq 2$. Then the following conditions are equivalent:
+    1. (i)
+      $e_{HK}(P)=\dfrac{8}{3}$.
+    2. (ii)
+      $\widehat{P}\cong k[[x_{1},\dots,x_{4},y_{1},\dots,y_{4}]]/(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2},y_{1}^{2}+y_{2}^{2}+y_{3}^{2}+y_{4}^{2},x_{i}y_{j})$.
+Proof.
+(1): Since $P$ is a three-dimensional unmixed local ring, by Proposition 5.4, both $R$ and $S$ must be three-dimensional and unmixed.
+Thus, since $R$ and $S$ are non-regular, by [31, Theorem 3.1(1)], we have $e_{HK}(R)\geq\dfrac{4}{3}$ and $e_{HK}(S)\geq\dfrac{4}{3}$.
+Combining this with Theorem 3.2, we obtain
+$e_{HK}(P)=e_{HK}(R)+e_{HK}(S)\geq\frac{8}{3}.$
+(2): (i) $\Rightarrow$ (ii) By hypothesis and Theorem 3.2, since $R$ and $S$ share the same dimension, we have:
+$e_{HK}(P)=e_{HK}(R)+e_{HK}(S)=\frac{8}{3}.$
+From [31, Theorem 3.1(1)], as $R$ and $S$ are non-regular, $e_{HK}(R)\geq\dfrac{4}{3}$ and $e_{HK}(S)\geq\dfrac{4}{3}$.
+Therefore, the equality $e_{HK}(R)+e_{HK}(S)=\dfrac{8}{3}$ implies that $e_{HK}(R)=\dfrac{4}{3}$ and $e_{HK}(S)=\dfrac{4}{3}$.
+Now, by [31, Theorem 3.1(2)], this yields
+$\widehat{R}\cong\frac{k[[x_{1},\dots,x_{4}]]}{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2})}\quad\text{and}\quad\widehat{S}\cong\frac{k[[y_{1},\dots,y_{4}]]}{(y_{1}^{2}+y_{2}^{2}+y_{3}^{2}+y_{4}^{2})}.$
+The desired result then follows from Fact 2.7.
+(i) $\Leftarrow$ (ii) Set $H=k[[x,y,z,w]]/(x^{2}+y^{2}+z^{2}+w^{2})$.
+We must then have that $\widehat{P}\cong H\times_{k}H$ (see Fact 2.7).
+Also, $e_{HK}(H)=\frac{4}{3}$ (see e.g., [31, Theorem 3.1(2)]).
+By the flatness of the canonical map $f:P\to\widehat{P}$, using Proposition 2.1(8) and Theorem 3.2, we have
+$e_{HK}(P)=e_{HK}(\widehat{P})=2e_{HK}(H)=\frac{8}{3}.$
+Let $(A,\mathfrak{m},k)$ be a local ring of characteristic $p>0$.
+We say that $A$ satisfies condition (#)(\#) if $\mathfrak{m}$ has a minimal reduction $J$.
+This condition will be considered in the next theorem.
+5.7 Theorem.
+Let $P=R\times_{k}S$ be a three-dimensional unmixed local ring of characteristic $p>0$.
+Assume further that $k=\overline{k}$ is algebraically closed, $p\neq 2$ and that both $R$ and $S$ are non-regular of characteristic $p$ satisfying the condition (#\#).
+Consider the following assertions:
+  1. (1)
+    $\frac{8}{3}<e_{HK}(P)\leq\frac{17}{6}$
+  2. (2)
+    Either $\widehat{P}\cong k[[x_{1},\dots,x_{4},y_{1},\dots,y_{4}]]/(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{c_{1}},y_{1}^{2}+y_{2}^{2}+y_{3}^{2}+y_{4}^{2},x_{i}y_{j})$ or $P\cong k[[x_{1},\dots,x_{4},y_{1},\dots,y_{4}]]/(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{c_{1}},y_{1}^{2}+y_{2}^{2}+y_{3}^{2}+y_{4}^{c_{2}},x_{i}y_{j})$ for integers $c_{1},c_{2}\geq 3$.
+  3. (3)
+    Either $\operatorname{gr}_{\mathfrak{m}}(P)\cong k[[x_{1},x_{2},x_{3},y_{1},\dots,y_{4}]]/(x_{1}^{2}+x_{2}^{2}+x_{3}^{2},y_{1}^{2}+y_{2}^{2}+y_{3}^{2}+y_{4}^{2},x_{i}y_{i})$ or $\operatorname{gr}_{\mathfrak{m}}(P)\cong k[x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}]/(x_{1}^{2}+x_{2}^{2}+x_{3}^{2},y_{1}^{2}+y_{2}^{2}+y_{3}^{2},x_{i}y_{j})$
+Then (1) $\Rightarrow$ (2) $\Rightarrow$ (3).
+Moreover, when (2) holds true, we have $e_{HK}(P)\geq\dfrac{17}{6}-\dfrac{2}{3c_{1}^{2}}$ or $e_{HK}(P)\geq 3-\dfrac{2}{3c_{1}^{2}}-\dfrac{2}{3c_{2}^{2}}$, respectively.
+Proof.
+(1) $\Rightarrow$ (2) Let $\frac{8}{3}<e_{HK}(P)\leq\frac{17}{6}$.
+Since $P=R\times_{k}S$ is a three-dimensional unmixed local ring, by Proposition 5.4, both $R$ and $S$ must be three-dimensional and unmixed.
+By [31, Theorem 3.1(1)], we have $e_{HK}(R)\geq\dfrac{4}{3}$ and $e_{HK}(S)\geq\dfrac{4}{3}$.
+Combining this with Theorem 3.2 and the inequalities $\frac{8}{3}<e_{HK}(P)\leq\frac{17}{6}$, without loss of generality, we may assume that $e_{HK}(S)=\frac{4}{3}$ and $\frac{4}{3}<e_{HK}(R)\leq\frac{3}{2}$ or $\frac{4}{3}<e_{HK}(R),e_{HK}(S)\leq\frac{3}{2}$.
+In the first case, we have
+(5.2) $R\cong\frac{k[[x_{1},\dots,x_{4}]]}{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{c_{1}})}\quad\text{and}\quad\widehat{S}\cong\frac{k[[y_{1},\dots,y_{4}]]}{(y_{1}^{2}+y_{2}^{2}+y_{3}^{2}+y_{4}^{2})},$
+by [31, Corollary 3.11] and [31, Theorem 3.1(2)], respectively. Therefore, the isomorphism for $\widehat{P}$ follows by Fact 2.7. In the second case, we have:
+$R\cong\frac{k[[x_{1},\dots,x_{4}]]}{(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{c_{1}})}\quad\text{and}\quad S\cong\frac{k[[y_{1},\dots,y_{4}]]}{(y_{1}^{2}+y_{2}^{2}+y_{3}^{2}+y_{4}^{c_{2}})},$
+by [31, Corollary 3.11] and the isomorphism for $P$ follows by Fact 2.7.
+[…]
+The ”moreover” part follows by Theorem 3.2, combining with Fact 2.7 and the formulas given in [31, Theorem 3.1] and [31, Corollary 3.11]. ∎
+5.8 Proposition.
+Let $R$ and $S$ be Cohen-Macaulay local rings of characteristic $p>0$, both of dimension 2.
+Let $T$ be a one-dimensional Cohen-Macaulay local ring of characteristic $p$, and suppose that the residue field $k$ is algebraically closed.
+Then $R\times_{T}S$ is not of ”minimal” Hilbert-Kunz multiplicity, that is, $e_{HK}(R\times_{T}S)>\frac{e(R\times_{T}S)+1}{2}$.
+In other words, the associated graded ring $\operatorname{gr}_{\mathfrak{m}}(R\times_{T}S)$ is not isomorphic to the Veronese subring $k[U,V]^{(e)}$, where $e=e(R)+e(S)$.
+Proof.
+First, observe that by [14, Lemma 2.1], the fiber product $R\times_{T}S$ is Cohen-Macaulay of dimension 2.
+Then, $e_{HK}(R\times_{T}S)\geq\frac{e(R\times_{T}S)+1}{2}$, by [30, Theorem 2.5].
+Suppose, for contradiction, that
+$e_{HK}(R\times_{T}S)=\frac{e(R\times_{T}S)+1}{2}$
+From this equality, by Theorem 3.2 and [18, Corollary 3.4], we derive
+$e_{HK}(R)+e_{HK}(S)=\frac{e(R)+e(S)+1}{2}.$
+Rewriting this, we get $(2e_{HK}(R)-e(R))+(2e_{HK}(S)-e(S))=1$.
+However, this leads to a contradiction as $2e_{HK}(R)-e(R)\geq 1$ and $2e_{HK}(S)-e(S)\geq 1$, by [30, Theorem 2.5] again.
+Then, we must have that $e_{HK}(R\times_{T}S)>\frac{e(R\times_{T}S)+1}{2}$.
+The ”in other words” part follows directly from the equivalence given in [30, Theorem 3.1]. ∎
+5.9 Example.
+Let $X\subset\mathbb{A}^{3}_{k}$ and $Y\subset\mathbb{A}^{3}_{k}$ be affine algebraic surfaces over an algebraically closed field $k$ of characteristic $p>0$, and let $Z=X\cap Y$.
+Suppose that the local rings at a point $x\in Z$,
+$\mathcal{O}_{X,x}\cong k[X]_{x}\quad\text{and}\quad\mathcal{O}_{Y,x}\cong k[Y]_{x},$
+are both Cohen-Macaulay of dimension 2, and that $T=\mathcal{O}_{Z,x}\cong k[Z]_{x}$ is Cohen–Macaulay of dimension 1.
+Then, the associated graded ring of the local ring of the union $X\cup Y$ at $x$ satisfies:
+$\operatorname{gr}_{\mathfrak{m}}(\mathcal{O}_{X\cup Y,x})\not\cong k[U,V]^{(e)}\subseteq k[U,V],$
+where $e=\operatorname{mult}_{x}(X)+\operatorname{mult}_{x}(Y)$, and $\operatorname{mult}_{x}(-)$ denotes the multiplicity of the corresponding variety at $x$ (i.e., the Hilbert-Samuel multiplicity of $\mathcal{O}_{X,x}$ or $\mathcal{O}_{Y,x}$).
+6. Bounds for the Hilbert-Kunz multiplicity of fiber products and idealizations
+In this section, we establish explicit lower bounds for the Hilbert-Kunz multiplicity of fiber products and idealization rings.
+These bounds are derived by combining our previous results with key known results from recent literature.
+Although we focus on specific classes of rings, our lower bounds provide significant improvements over existing bounds.
+Moreover, our results remain valid without certain restrictive assumptions that were crucial in previous results (see, for instance [1, 8, 7]).
+Note, for example, that the ring $R\times_{T}S$ is not necessarily unmixed, even if $R$ and $S$ are unmixed.
+Indeed, if $R=k[[x,y,z]]/(xy)$ and $S=k[[w]]$, we must then have that $R\times_{k}S\cong k[[x,y,z,w]]/(xy,xw,yw,zw),$ is not quasi-unmixed, as it has components of dimensions 1 and 2 (see Lemma 5.3).
+Remarkably, even without the unmixedness assumption required in Theorem 2.2, our results on fiber products yield substantially sharper lower bounds than those obtained in Theorem 2.2 and [8, Theorem 3.7].
+6.1 Theorem.
+Let $P=R\times_{T}S$ be a $d$-dimensional fiber product ring, where $R$ and $S$ are both (or at least one is) unmixed local rings of characteristic $p>0$ with $d:=\dim R\geq 2$. Then,
+$e_{HK}(R\times_{T}S)\Rightarrow\left\{\begin{array}[]{llll}=2;&\dim R=\dim S>\dim T(R,S\mbox{ regular})\\ \geq 2+\frac{1}{d\big{(}d!(d-1)+1\big{)}^{d}};&\dim R=\dim S>\dim T(R\mbox{ not-regular})\\ \geq 2\left(1+\frac{1}{d\big{(}d!(d-1)+1\big{)}^{d}}\right);&\dim R=\dim S>\dim T(R,S\mbox{ non-regular})\\ \geq 1+\frac{1}{d\big{(}d!(d-1)+1\big{)}^{d}};&\dim R>\dim S\geq\dim T(R\mbox{ non-regular}).\end{array}\right$.
+Proof.
+Following by Theorem 3.2 and Theorem $2.$ ∎
+6.2 Theorem.
+Let $P=R\times_{T}S$ be a local ring of characteristic $p>0$ with infinite perfect residue field $k$.
+Assume that $R$ and $S$ are unmixed local rings of dimension $d\geq 2$ and characteristic $p$, where both $R$ and $S$ are non-regular and $F$-finite.
+Let $(\boldsymbol{x})$ and $(\boldsymbol{y})$ be minimal reductions of $\mathfrak{m}_{R}$ and $\mathfrak{m}_{S}$ respectively, and define:
+  1. (1)
+    $\mu_{1}:=\max\{\mu(I)\mid I\text{ is an ideal of }R/(\boldsymbol{x})\}$ and $\mu_{2}:=\max\{\mu(I)\mid I\text{ is an ideal of }S/(\boldsymbol{y})\},$
+  2. (2)
+    $t_{1}:=\max\{n\in\mathbb{N}\mid\overline{\mathfrak{m}_{R}^{n}}\not\subseteq(\boldsymbol{x})\}$ and $t_{2}:=\max\{n\in\mathbb{N}\mid\overline{\mathfrak{m}_{S}^{n}}\not\subseteq(\boldsymbol{y})\}$.
+If $d>\dim T$, then
+$e_{HK}(P)\geq 2+\min\left\{\frac{1}{d!},\;\frac{\mu_{1}}{e(R)-\mu_{1}}\cdot\frac{1}{(\lceil d/t_{1}\rceil)^{d}}\right\}+\min\left\{\frac{1}{d!},\;\frac{\mu_{2}}{e(S)-\mu_{2}}\cdot\frac{1}{(\lceil d/t_{2}\rceil)^{d}}\right\},$
+Proof.
+The proof follows by Theorem 3.2 and [8, Theorem 3.7]. ∎
+Note that the fiber product $R\times_{T}S$ is not always Cohen-Macaulay or Gorenstein, even when $R$ and $S$ are Cohen-Macaulay or Gorenstein.
+For example, consider the rings $R$ and $S$ in Example 3.9, which are both 2-dimensional Cohen-Macaulay (in fact Gorenstein) hypersurfaces.
+Nevertheless, their fiber product $R\times_{k}S$ fails to be Cohen-Macaulay (and consequently not Gorenstein), as $\dim(R\times_{k}S)=2$, while we have $\text{depth}(R\times_{k}S)=1$.
+What is particularly noteworthy is that our results yield equally strong - and in fact sharper - lower bounds for both fiber products and idealizations, without requiring the Cohen-Macaulay or Gorenstein assumptions needed in [8, Theorem 1.2] and [7, Theorem 2.7].
+We will make these improved bounds explicit in the following results.
+.
+6.3 Theorem.
+Let $P=R\times_{T}S$ be a local ring of characteristic $p>0$ with infinite perfect residue field $k$.
+Let $R$ and $S$ be reduced $F$-finite Gorenstein rings of dimension $d>\dim T$ and characteristic $p$, and let $s_{1}$ and $s_{2}$ the $F$-signatures [^1] For the definition of F-signature, consult [20] of $R$ and $S$, respectively.
+Let $(\boldsymbol{x})$ and $(\boldsymbol{y})$ be minimal reductions of $\mathfrak{m}_{R}$ and $\mathfrak{m}_{S}$, respectively.
+For any ideals $I$ and $J$ of $R$ and $S$ (resp.) such that $(\boldsymbol{x})\subseteq I$ and $(\boldsymbol{y})\subseteq J$, let $\mu_{1}=\mu(I/(\boldsymbol{x}))$ and $\mu_{2}=\mu(J/(\boldsymbol{y}))$, then
+$e_{HK}(P)\geq 2+\frac{\mu_{1}}{e(R)-\mu_{1}}(1-s_{1})+\frac{\mu_{2}}{e(S)-\mu_{2}}(1-s_{2}),$
+Proof.
+Following by Theorem 3.2 and by [8, Theorem 1.2]. ∎
+6.4 Theorem.
+Let $P=R\times_{T}S$ be a local ring of characteristic $p>0$, and set $\alpha(-):=\max\left\{\frac{1}{p^{d}d!},\frac{1}{p^{d}e(-)}\right\}.$ Then,
+  • (1)
+    If $R$ and $S$ are non-regular Cohen-Macaulay rings of dimension $d>\dim T$, then
+    $e_{HK}(P)>2+\alpha(R)+\alpha(S).$
+  • (2)
+    If $R$ is regular and $S$ is a non-regular Cohen-Macaulay ring with $\dim R=\dim S=d>\dim T$, then
+    $e_{HK}(P)>2+\alpha(S).$
+  • (3)
+    If $R$ is a non-regular Cohen-Macaulay ring with $\dim R=d>\dim S\geq\dim T$, then
+    $e_{HK}(P)>1+\alpha(R).$
+Proof.
+Following by Theorem 3.2 and [7, Theorem 2.7]. ∎
+6.5 Remark.
+Let $A$ be a Noetherian local ring of characteristic $p>0$ with dimension $d\geq 2$ and let $M$ be a finitely generated $A$-module.
+Suppose that $\Lambda=\{\mathfrak{p}\in\operatorname{Assh}(A)\cap\operatorname{Supp}(M)\mid A/\mathfrak{p}\mbox{ is unmixed not regular}\}$.
+Then, by Associativite Formula (Proposition 2.1(5)) and by Theorem 2.2, we get
+$e_{HK}(\mathfrak{m},M)\geq\#(\Lambda)\left(1+\frac{1}{d\big{(}d!(d-1)+1\big{)}^{d}}\right).$
+In particular, if $M=A$ is unmixed and not regular, we get
+$e_{HK}(A)\geq\#(\Lambda)\left(1+\frac{1}{d\big{(}d!(d-1)+1\big{)}^{d}}\right).$
+This last formula improves the formula given by Theorem 2.2 for rings $A$ having at least two distinct primes $\mathfrak{p}\in\operatorname{Assh}(A)$ such that $A/\mathfrak{p}$ is unmixed and not regular.
+For example, if $P=R\times_{k}S$ with $\dim R=\dim S$, it has at least two distinct primes $\mathfrak{p}\oplus\mathfrak{m}_{S}$ and $\mathfrak{m}_{R}\oplus\mathfrak{q}$ in $\operatorname{Assh}(P)$, and it suffices to require that both $R/\mathfrak{p}$ and $S/\mathfrak{q}$ are unmixed and not regular.
+6.6 Proposition.
+Let $R$ be an unmixed local ring of characteristic $p>0$ and dimension $d\geq 2$. Suppose that $R$ is not regular, and let $M$ be a finitely generated $R$-module. Then,
+  • (1)
+    $e_{HK}(R\ltimes M)\geq(1+\#\Lambda)\left(1+\frac{1}{d\left(d!(d-1)+1\right)^{d}}\right).$
+  • (2)
+    If $R$ is a domain, $e_{HK}(R\ltimes M)\geq(\operatorname{rk}_{R}(M)+1)\left(1+\frac{1}{d\left(d!(d-1)+1\right)^{d}}\right).$
+Proof.
+(1) Follows from Corollary 4.3(1), combined with Theorem 2.2 and Remark 6.5, and (2) follows from Corollary 4.5(3) and Theorem $2.$ ∎
+6.1. Multi-factor fiber product rings
+We also consider multi-factor fiber products in the local Noetherian setting, which we now define.
+6.7 Definition.
+Let $R_{1},\dots,R_{r}$ and $T$ be Noetherian local rings and let $\pi_{i}:R_{i}\to T$ be surjective homomorphisms of local rings. We define the fiber product of $R_{1},\dots,R_{r}$ over $T$ as
+$R_{1}\times_{T}\cdots\times_{T}R_{r}=\{(a_{1},\dots,a_{r})\in R_{1}\oplus\cdots\oplus R_{r}\mid\pi_{i}(a_{i})=\pi_{j}(a_{j}),\ 1\leq i,j\leq r\}.$
+6.8 Remark.
+The multi-factor fiber product construction coincides with iteratively applying the two-factor fiber product construction to the sequence $R_{1},\dots,R_{r}$ a total of r1r-1 times, that is,
+$R_{1}\times_{T}\cdots\times_{T}R_{r}=((R_{1}\times_{T}R_{2})\times_{T}\cdots)\times_{T}R_{r}.$
+From Remark 6.8, we get that the multi-factor fiber product inherits almost all intrinsic properties from the two-factor fiber product of rings.
+In particular, properties such as Noetherianity, locality, and the behavior of the Krull dimension:
+(6.1) $\dim(R_{1}\times_{T}\cdots\times_{T}R_{r})=\max\{\dim R_{i}\mid 1\leq i\leq r\}\geq\min\{\dim R_{i}:1\leq i\leq r\}\geq\dim T,$
+as well as the property that $R_{1}\times_{T}\cdots\times_{T}R_{r}$ is a non-regular ring, among others, are preserved.
+From the (in)equalitie (6.1), it is easy to see that: If $d=\dim(R_{1}\times_{T}\cdots\times_{T}R_{r})=\dim T$, then $\dim R_{i}=d$ for $1\leq i\leq r$.
+Thus, the following formula for the Hilbert-Kunz multiplicity of the multi-factor fiber product $R_{1}\times_{T}\cdots\times_{T}R_{r}$ is obtained by applying Theorem 3.2 (r1)(r-1) times, as follows:
+6.9 Theorem.
+Let $P=R_{1}\times_{T}\cdots\times_{T}R_{r}$ be the multi-factor fiber product ring, and let $d=\dim(P)$. Let $R_{1},\dots,R_{r}$ and $T$ be Noetherian local rings of characteristic $p>0$. Then,
+$e_{HK}(P)=\begin{cases}\sum\limits_{\begin{subarray}{c}1\leq i\leq r\end{subarray}}e_{HK}(R_{i})-(r-1)e_{HK}(T),&d=\dim(T),\\[10.0pt] \sum\limits_{\begin{subarray}{c}1\leq i\leq r\\ \dim(R_{i})=d\end{subarray}}e_{HK}(R_{i}),&d>\dim(T).\end{cases}$
+Similarly to Theorem 6.1, we have the following generalization. This follows from Theorem 6.9 by applying Theorem 2.2 a finite number of times.
+6.10 Corollary.
+Let $P=R_{1}\times_{T}\cdots\times_{T}R_{r}$ be the multi-factor fiber product, and let $d=\dim(P)$.
+Let $R_{1},\dots,R_{r}$ and $T$ be Noetherian local rings of characteristic $p>0$ such that the rings $R_{i}$ of dimension $d$ are unmixed and non-regular.
+Then,
+$e_{HK}(P)\geq\begin{cases}r\left(1+\frac{1}{d\big{(}d!(d-1)+1\big{)}^{d}}\right)-(r-1)e_{HK}(T);&d=\dim(T),\\[10.0pt] t\left(1+\frac{1}{d\big{(}d!(d-1)+1\big{)}^{d}}\right);&d>\dim(T),\end{cases}$
+where $t$ denotes the number of indices $i$ for which $d=\dim R_{i}$, in case that $d>\dim T$.
+6.2. Application: Watanabe-Yoshida Conjecture for Fiber Products and Idealization Rings
+In this subsection, we present some immediate results regarding the Watanabe-Yoshida conjecture.
+Thanks to our results from previous sections, we show that there exists a large class of fiber product rings and Nagata idealizations which the conjecture holds under certain assumptions.
+We state the conjecture below.
 6.11 Conjecture (Watanabe-Yoshida [31]).
 Let $d\geq 2$ be an integer and $p>2$ be a prime number. Set
-$R_{p,d}:=\frac{k[[X_{0},X_{1},\ldots,X_{d}]]}{(X_{0}^{2}+\cdots+X_{d}^{2})}.$
-Let $(R,\mathfrak{m},k)$ be a $d$-dimensional unmixed local ring of characteristic $p$ and let $k=\overline{\mathbb{F}_{p}}$. Then the following statements hold:
+$A_{p,d}:=\frac{\overline{\mathbb{F}_{p}}[[X_{0},X_{1},\ldots,X_{d}]]}{(X_{0}^{2}+\cdots+X_{d}^{2})}.$
+Let $(A,\mathfrak{m},k)$ be a $d$-dimensional unmixed local ring of characteristic $p$ and let $k=\overline{\mathbb{F}_{p}}$. Then the following statements hold:
   1. (1)
-    If $R$ is not regular, then $e_{\text{HK}}(R)\geq e_{\text{HK}}(R_{p,d})\geq 1+m_{d}$.
+    If $A$ is not regular, then $e_{\text{HK}}(A)\geq e_{\text{HK}}(A_{p,d})\geq 1+m_{d}$.
   2. (2)
-    If $e_{\text{HK}}(R)=e_{\text{HK}}(R_{p,d})$, then the $\mathfrak{m}$-adic completion of $R$ is isomorphic to $R_{p,d}$ as local rings.
+    If $e_{\text{HK}}(A)=e_{\text{HK}}(A_{p,d})$, then the $\mathfrak{m}$-adic completion of $A$ is isomorphic to $A_{p,d}$ as local rings.
 Here, the numbers $m_{d}$ are defined by the relation
 $\sec x+\tan x=1+\sum_{d=1}^{\infty}m_{d}x^{d}\quad(-\pi/2<x<\pi/2).$
-Recently, Cox and Aberbach [8] proved that the Watanabe-Yoshida conjecture holds up to dimension $7$.
-The first inequality of part (1) of the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11].
-Also, recently, Trivedi [22] proved the second inequality of part (1) of the conjecture, for $d\geq 4$ and $p>d-3$.
-In our work, we calculated their Hilbert-Kunz multiplicity and gave conditions for fiber products and idealization rings not to be regular.
-The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas presented in the sections 3 and 4.
-5.4 Proposition.
-Let $d\geq 2$ be an integer and let $p>2$ be a prime number.
-Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p$ as in (2.5) and assume $R$ is a $d$-dimensional complete intersection, not regular.
-If $\dim R\geq\dim S>\dim T$, then
-$e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p}).$
-If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the fiber product $R\times_{T}S$.
+This conjecture has a positive answer for local rings with dimension $\leq 7$ (see [10]), but for larger dimensions, the problem remains open.
+Some other positive answers in certain cases are provided, for instance, in [28, 13].
+For the remainder of this section, we say that a $d$-dimensional Noetherian local ring $A$ of characteristic $p>0$ satisfies (WY1) provided that $A$ satisfies the strong inequality $e_{HK}(A)\geq e_{HK}(A_{p,d})$ of Watanabe-Yoshida conjecture 6.11(1).
+In our work, we obtained explicit formulas for the Hilbert-Kunz multiplicity of fiber products and idealization rings and proved that these rings are not regular.
+We use these formulas to obtain new classes of singular rings that satisfy (WY1).
+6.12 Theorem.
+Let $R$, $S$, and $T$ be Noetherian local rings of characteristic $p>0$. Assume that $\dim R=\dim R\times_{T}S$. If $R$ satisfies (WY1), then the fiber product $R\times_{T}S$ satisfies (WY1).
 Proof.
-The proof follows by Proposition 3.2, [11, Theorem 4.6] and [22, Theorem (A)]. ∎
+By the explicit formulas given in Theorem 3.2, it is easy to see that this assertion holds when $\dim R\geq\dim S>\dim T$.
+Thus, we may assume that $\dim R=\dim S=\dim T$, and then $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$.
+It is easy to see that $\pi_{S}(\mathfrak{m}_{S})=\mathfrak{m}_{T}$ and considering the exact sequence of $S$-modules
+$0\rightarrow\operatorname{Ker}\pi_{S}\rightarrow S\xrightarrow{\pi_{S}}T\rightarrow 0,$
+we have:
+$e_{HK}(S)-e_{HK}(\mathfrak{m}_{S},T)$ $=e_{HK}(\mathfrak{m}_{S},\operatorname{Ker}\pi_{S})$
+$e_{HK}(S)-\lim_{e\to\infty}\frac{\ell_{S}(T/\mathfrak{m}_{S}^{[q]}T)}{q^{\dim S}}$ $=e_{HK}(\mathfrak{m}_{S},\operatorname{Ker}\pi_{S})$
+$e_{HK}(S)-e_{HK}(T)=e_{HK}(S)-\lim_{e\to\infty}\frac{\ell_{T}(T/\mathfrak{m}_{T}^{[q]}T)}{q^{\dim T}}$ $=e_{HK}(\mathfrak{m}_{S},\operatorname{Ker}\pi_{S})\geq 0.$
+Thus, since $R$ satisfies (WY1), we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)\geq e_{HK}(A_{p,d})$. ∎
 6.13 Corollary.
-Let $d\geq 2$ be an integer, $p>2$ be a prime number, $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular, and let $I$ be an ideal of $R$.
-If $\dim R/I<\dim R$, then
-$e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p}).$
-If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the amalgameted duplication $R\bowtie I$.
-5.6 Proposition.
-Let $d\geq 2$ be an integer, $p>2$ be a prime number and let $(R,\mathfrak{m},k)$ be a $d$-dimensional complete intersection of characteristic $p$, not regular. If $M$ is a finite $R$-module, then
-$e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p}).$
-If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the idealization ring $R\ltimes M$.
+Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $I$ be a proper ideal of $R$. If $R$ satisfies (WY1), then the amalgamated duplication $R\bowtie I$ satisfies (WY1).
+6.14 Theorem.
+Let $R$ be a Noetherian local ring of characteristic $p>0$ and let $M$ be a finitely generated $R$-module. If $R$ satisfies (WY1), then the idealization $R\ltimes M$ satisfies (WY1).
 Proof.
-The proof follows by item (1) of Corollary 4.3, [11, Theorem 4.6] and [22, Theorem (A)]
-5.7 Remark.
-Propositions 5.2 and 5.4, 5.6 effectively expand the class of rings and ideals that satisfy the desired inequalities of conjectures 5.1 and 5.3, respectively, because an idealization ring $R\ltimes M$ and a fiber product ring $R\times_{T}S$ may not even be Cohen-Macaulay when $R$ is a complete intersection local ring.
-Both situations are exemplified below.
-5.8 Example ([3], Example 4.15).
-Let $(R,\mathfrak{m})$ be an $d$-dimensional regular local ring with $x_{1},\dots,x_{d}$ a minimal basis for $\mathfrak{m}$.
-Let $M_{i}=R/(x_{1},\dots,x_{i})$, $0\leq i\leq d$ (for $i=0$, $M_{i}=R$).
-Then $\dim R\ltimes M_{i}=d$ and $\operatorname{depth}R\ltimes M_{i}=d-i$.
-So $R\ltimes M_{i}$ is not Cohen-Macaulay, for $0<i\leq d$.
-5.9 Example.
-Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$. Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2].
+The assertion follows from the formula given in Corollary 4.3(1). ∎
+An interesting consequence of Theorems 6.12 and 6.14 and Corollary 6.13 is due to a result of Enescu and Shimomoto [13, Theorem 4.6], which shows that every complete intersection ring (not regular) of dimension $d\geq 2$ and characteristic $p>2$ satisfies (WY1).
+It is important to observe that if $R$ is a complete intersection, the rings $R\times_{T}S$ and $R\ltimes M$ are not always complete intersections (see Examples 3.9 and 4.6).
+Therefore, the next result provides new classes of non-regular local rings that satisfy the strong inequality of Conjecture 6.11.
+6.15 Corollary.
+Let $d\geq 2$ and $p>2$. If $R$ is a $d$-dimensional complete intersection ring of characteristic $p$, not regular, then the following assertions hold:
+  1. (1)
+    If $d=\dim R\times_{T}S$, where $S,T$ are Noetherian local rings of characteristic $p$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(A_{p,d})$.
+  2. (2)
+    If $I$ is a proper ideal of $R$, then $e_{HK}(R\bowtie I)\geq e_{HK}(A_{p,d})$.
+  3. (3)
+    If $M$ is a finitely generated $R$-module, then $e_{HK}(R\ltimes M)\geq e_{HK}(A_{p,d})$.
+Another interesting consequence of Theorems 6.12 and 6.14 and Corollary 6.13 is due to a result of Aberbach of Enescu [2, Section 4], which shows that the Watanabe-Yoshida Conjecture 6.11(1) holds if the ring $R$ has Hilbert-Samuel multiplicity $e(R)\leq 5$.
+It is important to observe that even if $R$ satisfies $e(R)\leq 5$, the rings $R\times_{T}S$ and $R\ltimes M$ may have Hilbert-Samuel multiplicity greater than five in several cases (e.g., if $\dim R=\dim S>\dim T$ and $\dim R=\dim M$, respectively).
+Therefore, the next result provides new classes of non-regular local rings that satisfy the strong inequality of Conjecture 6.11.
+6.16 Corollary.
+Let $d\geq 2$ and $p>2$. If $R$ is a $d$-dimensional Noetherian local ring of characteristic $p$, not regular, with $e(R)\leq 5$, then the following assertions hold:
+  1. (1)
+    If $d=\dim R\times_{T}S$, where $S,T$ are Noetherian local rings of characteristic $p$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(A_{p,d})$.
+  2. (2)
+    If $I$ is a proper ideal of $R$, then $e_{HK}(R\bowtie I)\geq e_{HK}(A_{p,d})$.
+  3. (3)
+    If $M$ is a finitely generated $R$-module, then $e_{HK}(R\ltimes M)\geq e_{HK}(A_{p,d})$.
+Acknowlegment.
+The first author was supported by Sao Paulo Research Foundation (FAPESP) under grant 2019/21181-0. The second author was supported by Sao Paulo Research Foundation (FAPESP) under grant 2022/12114-0.
 Competing interests: The authors declare none.